### v1

Abstract
In this paper, we study the geometric nonlinearity properties, such as curvature and torsion, in a dual coordinate system of the Riemannian manifold defined by the Gaussian distribution. We also give the Amari formulas explicitly in this new coordinate system, which allows us to characterize existing geometric invariants, such as the dual potential function and the Fisher metric.
keywords:
Tensor curvature, Riemannian metric, Torsion.
1 Introduction
Information geometry, although a very well-known and rich branch of mathematics, has very important tools for studying certain geometric properties such as curvature and torsion, and for studying probability and information through differential geometry for example see [1]. The main idea is to study nonlinear geometric properties, such as curvature and torsion, in a dual coordinate system on a statistical manifold defined by a Gaussian distribution. Several researchers have studied this field, including S. I. Amari [2], one of the pioneers of information geometry. M. N. Boyom [3], who studies umerical properties of Koszul connections’. Koszul and Jean-Louis [4], who study deformations of locally flat connections. In [5], the geometric properties of beta distributions are studied. Calculating curvature and torsion is important because it allows a statistical manifold to be described globally. Curvature is used in fields such as physics. Since this manifold is torsion-free and curvature-free in the local coordinate system, can these properties be preserved in the dual coordinate system to conserve the obtained values? Research in this area dates back to Rao, who first recognized the importance of the differential geometry approach in statistics. He defined the Riemannian metric in terms of the Fisher information matrix, $g^{ij}$, and calculated the geodesic distances between two distributions for various statistical models. Amari [2], studies the geometric properties of the Riemannian manifold of a statistical model. He has shown that the curvature is entirely contained within the plane, and that the curvature in the $\theta$-coordinate system is constant. Building on Amari’s work, Ovidiu Calin[6] also presents the properties of the Gaussian model as a Riemannian manifold and its characteristics in the $\theta$-coordinate system. In [7], it then introduces the Riemann curvature tensor, and quickly moves on to sub-manifold theory in order to give the curvature tensor a concrete quantitative interpretation. In [8], it is shown that the statistical manifold define by lognormal distribution is flat with respect to a pair of torsion-free dual affine connections $\nabla$ and $\nabla^{*}$. Recently it has been observed that the space of fields $T^{l}(\theta)$ is seen as a subspace of the space of second degree trinomial with coefficients in $\sigma^{-3}C^{\infty}(S)$ where $\sigma^{-3}C^{\infty}(S)$ is a modulus $\mathbb{R}$. where $\sigma^{-3}C^{\infty}(S)$ is a modulus $\mathbb{R}$. In the same order as Amari and Bovai Lobaschevki on the manifold of models we have proved that the curvature in the $\xi$-coordinate system is equal to
$\mathcal{K}$ $=$ $\frac{16\mu^{2}\sigma^{8}+12\mu^{4}+6\sigma^{10}-\mu\sigma^{8}-2\mu^{2}\sigma^{2}+10\mu^{3}\sigma^{6}+92\mu^{3}\sigma^{8}+48\mu^{5}\sigma^{6}}{4\sigma^{10}}$
$+\frac{12\mu\sigma^{10}+24\mu^{5}-6\mu^{3}}{4\sigma^{10}}$
we will study the non-linear properties of the manifold in system of the $\xi$- coordinates which contains a gaussian model. An extension of this work would then be to formalize to what degree certain geometric information does not vary for the normal law whatever the reference field. After the introduction, the first section recall the preliminaries motion on theory of statistical manifold, in section $2$ we determine the Riemannian structure in $\xi$-coordinates system, in section $3$, we determine torsion tensor and curvature tensor. At the end, we present the conclusion.
2 Preliminaries
Soit $S=\left\{p_{\theta}(x),\left.\begin{array}[]{ll}\theta\in\Theta&\hbox{}\\ x\in\mathcal{X}&\hbox{}\end{array}\right.\right\}$ be the set of probabilities $p_{\theta}$, parameterized by $\Theta$, open subset of $\mathbb{R}^{n}$; on the sample space $\mathcal{X}\subseteq\mathbb{R}$, with $\theta=\left(\theta_{1},\theta_{2},\dots,\theta_{n}\right)$. In [8], the Gaussian model is a family of probabilistic distributions. consisting of density functions that are defined as follows $S=\left\{p_{\theta}(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}},\left.\begin{array}[]{ll}\theta\in\mathbb{R}\times\mathbb{R}^{*}_{+}&\hbox{}\\ x\in\mathbb{R}^{*}_{+}&\hbox{}\end{array}\right.\right\}$ with $p_{\theta}(x)$ the gaussian distribution parameterized by $\theta=(\mu,\sigma)$.
2.1 $\theta$-coordinates System and $\xi$-coordinates system..
Let $S$, the statistical manifold
Definition 1.
[2] Let $U$ be an open subset of $S$ and $\phi$ be a homeomorphism of $U$ in $\mathbb{R}^{n}$. Then a point $p\in U$ is a function de $\theta=\left(\theta_{1},\theta_{2},\dots,\theta_{n}\right)\in\mathbb{R}^{n}$ by $\phi$, i.e. $\phi(p)=\theta$ this application $\phi$ is called a function coordinates in the neighbourhood of $p$. It is introduced a coordinate system in $U$ called the Any point $p$ in $U$ is given by the $\theta$-coordinates given by $\theta=(\theta_{i});i\in\{1,2\dots,n\}$.
According to Amari’s[2], the $l-$representation basis is given by $\mathcal{B}_{\theta}l(x,\theta)=\left\{\partial_{1}l(x,\theta),\partial_{2}l(x,\theta)\right\}$ where $\mathcal{B}_{\theta}=\partial_{1},\partial_{2}\}$ is the natural basis. In the case of gaussian model we have $\partial_{1}=\frac{\partial}{\partial\mu}$, and $\partial_{2}=\frac{\partial}{\partial\sigma}$ therefore   $\left\{\frac{\partial}{\partial\mu},\frac{\partial}{\partial\sigma}\right\}=\mathcal{B}_{\theta}$ The base associated to the $l$- representation is therefore $\left\{\partial_{1}l(x,\theta),\partial_{2}l(x,\theta)\right\}$. As
$l(x,\theta)$ $=$ $-\log\sqrt{2\pi}\sigma-\frac{(x-\mu)^{2}}{2\sigma^{2}}$ (1)
It follows that
$\partial_{1}l(x,\theta)$ $=$ $-\frac{1}{\sigma}+\frac{(x-\mu)^{2}}{\sigma^{3}}$
So
$\mathcal{B}_{\theta}l(x,\theta)$ $=$ $\left\{\frac{(x-\mu)}{\sigma^{2}},-\frac{1}{\sigma}+\frac{(x-\mu)^{2}}{\sigma^{3}}\right\}.$
Remark 1.
[2] There are several other coordinate functions which at a point $p$, are such that: $\psi(p)=\xi=\left(\xi_{1},\xi_{2},\dots,\xi_{n}\right)$ where $\psi$ is a homeomorphism of an open $V$ of $S$ in $\mathbb{R}^{n}$ is the $\xi$-coordinate system. The coordinates $\left(\xi_{i}\right),\;i\in\{1,\dots,n\}$, characterize the point $p$ in $\mathbb{R}^{n}$. Let’s give ourselves two system coordinates $\theta$ and $\xi$ There is one and only one correspondence between the coordinates $\theta$ and $\xi$; given the properties enjoyed by the applications $\phi$ and   $\psi$, and the fact that they are homeomorphisms we have the following equations
$\xi$ $=$ $\psi\circ\phi^{-1}(\theta),\textrm{and}\;\theta=\phi\circ\psi^{-1}(\xi)$
The transformation of $\theta$ a would be a diffeomorphism, if the $n$-functions $\xi_{i}\left(\theta_{1},\dots,\theta_{n}\right)$ are differentiable and $Det\left|\frac{\partial\xi_{i}}{\partial\theta_{j}}\right|$ is non-zero.
2.2 Determination of the characteristic elements of the $\xi$- coordinates.
Let $S=\left\{p_{\theta}(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}},\left.\begin{array}[]{ll}\theta\in\mathbb{R}\times\mathbb{R}^{*}_{+}&\hbox{}\\ x\in\mathbb{R}^{*}_{+}&\hbox{}\end{array}\right.\right\}$ the gaussian statistical model. The tangent space of $l-$representation is given by
$T^{l}(\theta)$ $=$ =$=\left\{A(x);A(x)=A^{1}\frac{\partial_{1}l(x,\theta)}{\partial\mu}+A^{2}\frac{\partial_{2}l(x,\theta)}{\partial\sigma}\right\}$
and using (2.1) we have
$T^{l}(\theta)$ $=$ $\left\{A(x)\textrm{ tel que }A(x)=A^{1}_{(\mu;\sigma)}(x)\frac{x-\mu}{\sigma^{2}}+A^{2}_{(\mu;\sigma)}(x)\frac{(x-\mu)^{2}}{\sigma^{3}}-\frac{A^{2}_{(\mu;\sigma)}(x)}{\sigma}\right\};A^{1},A^{2}\in C^{\infty}(S)$
So $A(x)\in T^{l}(\theta)$ if and only if there exists $A^{1},A^{2}\in C^{infty}(S)$ such that
$A(x)$ $=$ $\frac{A^{2}}{\sigma^{3}}x^{2}+\left(\frac{A^{1}}{\sigma^{2}}-2\frac{A^{2}\mu}{\sigma^{3}}\right)x+\frac{A^{2}}{\sigma^{3}}\mu^{2}-\frac{A^{1}\mu}{\sigma^{2}}-\frac{A^{2}}{\sigma}$
Let $a=\frac{A^{2}}{\sigma^{3}}$ ; $b=\frac{A^{1}}{\sigma^{2}}-2\frac{A^{2}\mu}{\sigma^{3}}$ et $c=\frac{A^{2}}{\sigma^{3}}\mu^{2}-\frac{A^{1}\mu}{\sigma^{2}}-\frac{A^{2}}{\sigma}$ So $A(x)$ is of the form $ax^{2}+bx+c$ we can be reassured that $a=\frac{A^{2}}{\sigma^{3}}\neq 0$. We have thus prove the following result. Put $frac{1}{\sigma^{3}}C^{\infty}(S)_{2}[x]$ the modulus of polynomials of second degree with coefficients in $\frac{1}{\sigma^{3}}C^{\infty}(S)$. Then $T^{l}(\theta)$ is a subset of $\frac{1}{\sigma^{3}}C^{\infty}(S)_{2}[x]$.
Remark 2.
According to [2], it is shown that
$\mathbb{E}[\partial_{i}l(x,\theta)]$ $=$ $0$ (2)
Furthermore, it follows from that
$\mathbb{E}[A(x)]$ $=$ $A^{1}\mathbb{E}[\partial_{1}l(x,\theta)]+A^{2}\mathbb{E}[\partial_{2}l(x,\theta)].$
using (2) we obtain
$\mathbb{E}[A(x)]$ $=$ 00 0
[…]
$c$ $=$ $-a(\mu^{2}+\sigma^{2})-b\mu$
By setting $\xi_{2}=\mu^{2}+\sigma^{2}$ et $\xi_{1}=\mu$. $\xi=(\xi_{1},\xi_{2})=(\mu,\mu^{2}+\sigma^{2})$.
So $\xi=(\xi_{1},\xi_{2})=(\mu,\mu^{2}+\sigma^{2})$ is a coordinate system parameterized by $(\mu,\sigma)$. For it suffices show that $\theta$ is a diffeomorphism. Consider the application
$\theta:\mathbb{R}^{2}$$\rightarrow$ $\mathbb{R}^{2}:(\mu,\sigma)\mapsto(\theta_{1}(\mu,\sigma),\theta_{2}(\mu,\sigma))=(\xi_{1},\xi_{2})=(\mu,\mu^{2}+\sigma^{2})$
by noting that the components $\theta_{1}(\mu,\sigma)$ and $\theta_{2}(\mu,\sigma)$ are differentiable, it is sufficient to show that the Jacobian of $\theta$ is non-zero. Let $J_{M}=\left(\begin{array}[]{cc}\frac{\partial\theta_{1}}{\partial\mu}&\frac{\partial\theta_{1}}{\partial\sigma}\\ \frac{\partial\theta_{2}}{\partial\mu}&\frac{\partial\sigma_{2}}{\partial\sigma}\\ \end{array}\right)$
The Jacobian matrix of $\theta$. It follows that $J_{M}=\left(\begin{array}[]{cc}1&0\\ 2\mu&2\sigma\\ \end{array}\right)$ and its determinant which is the Jacobian is $2\sigma\neq 0$ because $\sigma>0$. It follows that according to the local inversion theorem is locally a diffeomorphism of class $C^{1}$. In coordinates we will need the natural basis and in the system of $\xi$- coordinates we will need the basis associated with the $l$- representation to be determined. To do this we use the Jacobian matrix of the basis passage $\{\mathcal{B}_{\theta}\}$ to the $\xi$-coordinate basis denoted by $\{\mathcal{B}_{\xi}\}=\{\partial_{\xi_{1}},\partial_{\xi_{2}}\}$. Denote by
(¯$(\overline{\mathcal{B}}^{i}_{\alpha})=\left(\begin{array}[]{cc}\frac{\partial\theta_{1}}{\partial\xi_{1}}&\frac{\partial\theta_{1}}{\partial\xi_{2}}\\ \frac{\partial\theta_{2}}{\partial\xi_{1}}&\frac{\partial\theta_{2}}{\partial\xi_{2}}\\ \end{array}\right)=J_{M}^{-1}$
the inverse of the Jacobian matrix $J_{M}$, and that
$(\partial_{i^{\prime}})=(\overline{\mathcal{B}}^{i}_{\alpha})(\partial_{i})$
. Furthermore, the inverse matrix
$J_{M}^{-1}=\left(\begin{array}[]{cc}1&0\\ \frac{-\mu}{\sigma}&\frac{1}{2\sigma}\\ \end{array}\right)$
So, the $\xi$- coordinates system is given by $\xi=(\xi_{1},\xi_{2})=(\mu,\mu^{2}+\sigma^{2})$, and we have the $l$- representation in the $\xi$-coordinates system is given by
$\mathcal{B}_{\xi}l(x,\xi)=\{\partial_{\xi_{1}}l(x,\xi),\partial_{\xi_{2}}l(x,\xi)\}$
where $\mathcal{B}_{\xi}=\{\partial_{\xi_{1}},\partial_{\xi_{2}}\}$ is the natural basis of the Amari associated manifold. where $\partial_{\xi_{1}}=\frac{\partial}{\partial\xi_{1}}$ and $\partial_{\xi_{2}}=\frac{\partial}{\partial\xi_{2}}$ thus  $\left\{\frac{\partial}{\partial\xi_{1}},\frac{\partial}{\partial\xi_{2}}\right\}=\mathcal{B}_{\xi}.$
Theorem 1.
Let $(S,g)$ the Riemannian manifold. The natural basis associated to $l$-representation in the coordinated system $(\xi_{1},\xi_{2})$ is given by
$\mathcal{B}(l,\xi)=\left\{\frac{x-\mu}{\sigma^{2}};\frac{(x-\mu)^{2}}{2\sigma^{4}}-\frac{\mu(x-\mu)}{\sigma^{3}}-\frac{1}{2\sigma^{2}}\right\}.$
Proof.
$(\partial_{\xi})=(\bar{\mathcal{B}}^{i}_{\alpha})(\partial_{i})$, $\alpha,i\in\left\{1,2\right\}$, we have
$\left(\begin{array}[]{c}\partial_{\xi_{1}}\\ \partial_{\xi_{2}}\\ \end{array}\right)=\left(\begin{array}[]{cc}1&0\\ -\frac{\mu}{\sigma}&\frac{1}{2\sigma}\\ \end{array}\right)\left(\begin{array}[]{c}\partial_{1}\\ \partial_{2}\\ \end{array}\right)$
So we have $\{\partial_{\xi_{1}},\partial_{\xi_{2}}\}=\{\partial_{1},\frac{-\mu}{\sigma}\partial_{1}+\frac{1}{2\sigma}\partial_{2}\}$ in $\theta$- coordinates. Its $l$-representation is $\partial_{\xi}l(x,\theta)\}=\left\{\partial_{1}l(x,\theta),-\frac{\mu}{\sigma}\partial_{1}l(x,\theta)+\frac{1}{2\sigma}\partial_{2}l(x,\theta)\right\}$
We have $\left\{\begin{array}[]{ll}\partial_{\xi_{1}}l(x,\theta)=\frac{x-\mu}{\sigma^{2}}&\hbox{}\\ \partial_{\xi_{2}}l(x,\theta)=\frac{(x-\mu)^{2}}{2\sigma^{4}}-\frac{\mu(x-\mu)}{\sigma^{3}}-\frac{1}{2\sigma^{2}}&\hbox{}\end{array}\right.$
Remark 3.
According to [2], the riemannian metric is given by
$g_{ij}(\theta)=<\partial_{i}l(x,\theta),\partial_{j}l(x,\theta)>\textrm{where},g_{ij}(\theta)=\mathbb{E}(\partial_{i}l(x,\theta).\partial_{j}l(x,\theta))=-\mathbb{E}(\partial_{i}\partial_{j}l(x,\theta)).$ (3)
and $g^{ij}$ will be its reverse called stamp information of Fisher. In the same, the metric one associated in the system from the $\xi$- coordinates is defined by
$g_{\alpha\beta}(\xi)=<\partial_{\xi_{\alpha}},\partial_{\xi_{\beta}}>;\;\alpha,\beta\in\{1,2\}$ (4)
2.3 Affine connection.
The collection of all $T_{p}S$ defined a $2n-$dimensional manifold $T(S)$ called Tangent bundle of $S$. More explicitly, $T(S)$ is completely defined by the projection mapping $\pi:TS\longrightarrow S$ such that $\pi^{-1}(\{p\})\cong\mathbb{R}^{n}$ and $\pi$ have the usual local trivialization property. $A:S\longrightarrow TS$ is a vector field on $S$ if $\pi\circ A=1_{S}$. The collection $\mathfrak{X}(S)$ of all vector fields on $S$ is an $C^{\infty}(S)-$module. An affine connection is an $\mathbb{R}$-bilinear mapping
$\nabla:\mathfrak{X}(S)\longrightarrow\mathfrak{X}(S):(A,B)\mapsto\nabla_{A}B$ such that:
$\nabla_{A}(fB)=(Af)(B)+f\nabla_{A}(B),\;\nabla_{fA}(B)=f\nabla_{A}(B).$
In local coordinate $\theta=\left(\theta_{1},\dots,\theta_{n}\right)$ of $S$ , $\nabla_{\partial_{i}}(\partial_{j})=\sum\Gamma_{ij}^{k}\partial_{i}$ and $\Gamma_{ij}^{k}$ is Christoffel coefficient given by: $\Gamma_{ij}^{k}=\frac{1}{2}g^{km}[\partial_{i}.g_{jm}+\partial_{j}.g_{ik}-\partial_{k}.g_{ij}].$
2.4 Torsion tensor and curvature of affine connection.
Let $\nabla$ be an affine-connection on $S$. A torsion tensor of $\nabla$ is an $\mathbb{R}$-bilinear mapping
$T:\mathfrak{X}(S)\times\mathfrak{X}(S)$$\longrightarrow$ $\mathfrak{X}(S):(A,B)\longmapsto T(A,B)=\nabla_{A}(B)-\nabla_{B}(A)-[A,B]$
In the case of statistical manifold $S=\left\{p_{\theta}(x),\left.\begin{array}[]{ll}\theta\in\Theta&\hbox{}\\ x\in\mathcal{X}&\hbox{}\end{array}\right.\right\}$. Amari [2], show that
$\mathbb{E}[\partial_{i}l(x,\theta)]$ $=$ $0$ (5)
Where $l(x,\theta)$ is the $l-$representation of $T_{\theta}$. In this case by taking into consideration the Amari [2], the result affine connection of $T_{\theta}$ on statistical manifold $S=\left\{p_{\theta}(x),\left.\begin{array}[]{ll}\theta\in\Theta&\hbox{}\\ x\in\mathcal{X}&\hbox{}\end{array}\right.\right\}$ is define by
$\Gamma_{ijk}:=\mathbb{E}\left[\partial_{i}\partial_{j}l(x,\theta).\partial_{k}l(x,\theta)\right].$ (6)
The family $\Gamma_{ijk}$ are called $l-$representation of coefficients of the connection $\nabla$ on $S$. With respect to the $l-$representation, $T_{ijk}(\theta)=\langle T(\partial_{\theta_{i}},\partial_{j}),\partial_{k}\rangle,$ it is the coefficient of $T$ with respect to $\partial_{\theta}l(x,\theta)$. Since $T(\partial_{i},\partial_{j})=\nabla_{\partial_{i}}(\partial_{j})-\nabla_{\partial_{j}}(\partial_{i})-(\partial_{i}\partial_{j}-\partial_{j}\partial_{i})$, then $T_{ijk}(\theta)=<\nabla_{\partial_{i}}(\partial_{j})-\nabla_{\partial_{j}}(\partial_{i})-(\partial_{i}\partial_{j}-\partial_{j}\partial_{i}),\partial_{k}>.$ We obtain the following expression: $T_{ijk}(\theta)=<\nabla_{\partial_{i}}(\partial_{j}),\partial_{k}>-<\nabla_{\partial_{j}}(\partial_{i}),\partial_{k}>.$ Therefore,
$T_{ijk}(\theta)=\Gamma_{ijk}(\theta)-\Gamma_{jik}(\theta).$ (7)
The curvature tensor of affine connection $\nabla$ on a $n$-manifold $S$ is a $\mathbb{R}-$trilinear mapping $R:\mathfrak{X}(S)\times\mathfrak{X}(S)\times\mathfrak{X}(S)\longrightarrow\mathfrak{X}(S):(X,Y,Z)\longmapsto R(X,Y)Z$
such that $R(X,Y)Z=[\nabla_{X},\nabla_{Y}]Z-\nabla_{[X,Y]}Z,$ with $[\nabla_{X},\nabla_{Y}]Z=\nabla_{X}\nabla_{Y}Z-\nabla_{Y}\nabla_{X}Z.$ In a locale coordinates $\theta=\left(\theta_{1},\dots,\theta_{n}\right)$ of $S$, we have $R(\partial_{i},\partial_{j}).\partial_{k}=\sum_{m=1}^{n}R_{ijk}^{m}\partial_{m}$
where $R_{ijk}^{m}=\frac{\partial\Gamma_{kl}^{m}}{\partial_{j}}-\frac{\partial\Gamma_{jl}^{m}}{\partial_{k}}+\Gamma_{js}^{m}\Gamma_{kl}^{s}-\Gamma_{ks}^{m}\Gamma_{jl}^{s}.$ Taking in consideration the $l-$representation of $T_{\theta}$ on statistical manifold $S=\left\{p_{\theta}(x),\left.\begin{array}[]{ll}\theta\in\Theta&\hbox{}\\ x\in\mathcal{X}&\hbox{}\end{array}\right.\right\}$. The family
$R_{ijkm}:=\langle R(\partial_{i},\partial_{j},\partial_{k}),\partial_{m}\rangle$ (8)
is the $l-$ representation of Riemann-Christoffel curvature tensor of the affine connection $\nabla$ on $S.$ It follow From (8) that $R_{ijkm}$ the following expression:
$R_{ijkm}=\left(\partial_{i}\Gamma^{s}_{jk}-\partial_{j}\Gamma^{s}_{ik}\right)g_{sm}+\left(\Gamma_{irm}\Gamma^{r}_{jk}-\Gamma_{jrm}\Gamma^{r}_{ik}\right).$ (9)
where $g_{sm}$ are coefficients of the Riemannian metric $g$ on $S$. A space with an affine connection is said to be flat when the Riemann-Christoffel curvature and torsion vanishes identically, i.e; $R_{ijkm}(\theta)=0$ holds for any $\theta$. In the coordinate system $\theta=\left(\theta_{1},\dots,\theta_{n}\right)$, the scalar curvature $\mathcal{K}$ of affine connection $\nabla$ is defined by
$\mathcal{K}=\frac{1}{n(n-1)}R_{ijkm}g^{im}g^{jk}.$ (10)
So we have the scalar curve in the frame of reference $\xi$ the noted number $\mathcal{K}$ define by
$\mathcal{K}=\frac{1}{2}R_{\alpha\beta\gamma\eta}g^{\alpha\eta}g^{\beta\gamma}.$ (11)
According to [2], and using 3, we have
$g_{11}(\theta)$ $=$ $-\mathbb{E}\left[\partial_{1}\partial_{1}l(x,\theta)\right]$
$=$ $-\mathbb{E}\left[-\frac{1}{\sigma^{2}}\right]$
$=$ $\frac{1}{\sigma^{2}}.$
$g_{22}(\theta)$ $=$ $-\mathbb{E}\left[\partial_{2}\partial_{2}l(x,\theta)\right]$
$=$ $-\mathbb{E}\left[\frac{1}{\sigma^{2}}-\frac{3(x-\mu)^{2}}{\sigma^{4}}\right]$
$=$ $-\mathbb{E}\left(\frac{1}{\sigma^{2}}\right)-\mathbb{E}\left[-\frac{3(x-\mu)^{2}}{\sigma^{4}}\right]$
$=$ $-\frac{1}{\sigma^{2}}+3\mathbb{E}\left[\frac{(x-\mu)^{2}}{\sigma^{4}}\right]$
$=$ $-\frac{1}{\sigma^{2}}+3\mathbb{E}\left[\partial_{1}l(x,\theta).\partial_{1}l(x,\theta)\right]$
$=$ $-\frac{1}{\sigma^{2}}-3\mathbb{E}\left[\partial_{1}\partial_{1}l(x,\theta)\right]$
$=$ $-\frac{1}{\sigma^{2}}-3\mathbb{E}\left[-\frac{1}{\sigma^{2}}\right]$
$=$ $-\frac{1}{\sigma^{2}}+\frac{3}{\sigma^{2}}$
$=$ $\frac{2}{\sigma^{2}}$
$g_{12}(\theta)$ $=$ $-\mathbb{E}\left[\partial_{1}\partial_{2}l(x,\theta)\right]$
$=$ $-\mathbb{E}\left[\frac{2}{\sigma^{3}}(x-\mu)\right]$
$=$ $-\frac{2}{\sigma^{\sigma}}\mathbb{E}\left[\frac{(x-\mu)}{\sigma^{2}}\right]$
$=$ $-\frac{2}{\sigma}\mathbb{E}\left[\partial_{1}l(x,\theta)\right]$
$=$ 00 0
$g_{21}(\theta)$ $=$ $-\mathbb{E}\left[\partial_{2}\partial_{1}l(x,\theta)\right]$
$=$ $-\mathbb{E}\left[-\frac{2}{\sigma^{3}}(x-\mu)\right]$
$=$ $\frac{2}{\sigma^{3}}\mathbb{E}\left[(x-\mu)\right]$
$=$ $\frac{2}{\sigma}\mathbb{E}\left[\frac{(x-\mu)}{\sigma^{2}}\right]$
$=$ $\frac{2}{\sigma}\mathbb{E}\left[\partial_{1}l(x,\theta)\right]$
$=$ 00 0
So, the metric one associated for a manifold of gaussian model $S$ is given by
$g_{ij}(\theta)=\left(\begin{array}[]{cc}\frac{1}{\sigma^{2}}&0\\ 0&\frac{2}{\sigma^{2}}\\ \end{array}\right)$ (12)
3 Riemannian metric in $\xi$-coordinates system.
Theorem 2.
Let $\xi=\left(\mu,\mu^{2}+\sigma^{2}\right)$ the $\xi$-coordinate system on gaussian manifold. Let $\left(g_{ij}\right)_{1\leq i,j\leq 2}$ the Riemannian of metric in $\theta-$ coordinate on $S$ given by equation (12), and (¯$(\overline{\mathcal{B}}^{i}_{\alpha})=\left(\begin{array}[]{cc}1&0\\ \frac{-\mu}{\sigma}&\frac{1}{2\sigma}\\ \end{array}\right)$ The inverse of the jacobian matrix. The Riemannian of metric $g_{\alpha\beta}(\xi)$, in the $\xi$- co-ordinates is given by
$g_{\alpha\beta}(\xi)=\bar{B_{\alpha}}^{1}\bar{B_{\beta}}^{1}g_{11}(\theta)+\bar{B_{\alpha}}^{2}\bar{B_{\beta}}^{2}g_{22}(\theta)$
$\alpha,\;\beta\in\left\{1,2\right\}$.
Proof.
$g_{\alpha\beta}(\xi)=<\partial_{\xi_{\alpha}},\partial_{\xi_{\beta}}>,\;\alpha,\;\beta\in\left\{1,2\right\}.$
or $\bar{B_{\alpha}}^{i}=\left(\frac{\partial\theta_{i}}{\partial\xi_{\alpha}}\right)$ with
$\partial_{\xi_{\alpha}}=\sum\limits_{i=1}^{2}\bar{B_{\alpha}}^{i}\partial_{i}=B_{\alpha}^{1}\partial_{1}+B_{\alpha}^{2}\partial_{2},\;and\;\partial_{\xi_{\beta}}=\sum\limits_{j=1}^{2}\bar{B_{\beta}}^{j}\partial_{j}=B_{\beta}^{1}\partial_{1}+B_{\beta}^{2}\partial_{2}.$
We have
$g_{\alpha\beta}(\xi)$ $=$ $<B_{\alpha}^{1}\partial_{1}+B_{\alpha}^{2}\partial_{2},B_{\beta}^{1}\partial_{1}+B_{\beta}^{2}\partial_{2}>$
$=$ $<B_{\alpha}^{1}\partial_{1},B_{\alpha}^{1}\partial_{1}>+<B_{\alpha}^{1}\partial_{1},B_{\alpha}^{2}\partial_{2}>+<B_{\alpha}^{2}\partial_{2},B_{\alpha}^{1}\partial_{1}>+<B_{\alpha}^{2}\partial_{2},B_{\alpha}^{2}\partial_{2}>$
$=$ $\bar{B_{\alpha}}^{1}\bar{B_{\beta}}^{1}g_{11}(\theta)+\bar{B_{\alpha}}^{2}\bar{B_{\beta}}^{2}g_{22}(\theta)$
$\forall\alpha,\;\beta\in\left\{1,2\right\}$. ∎
While being (12) and (2) there is the proof of following theorem
Theorem 3.
Let $\xi=\left(\mu,\mu^{2}+\sigma^{2}\right)$ the dual coordinate system of $\theta-$ coordinate system on gaussian manifold. The riemannian metric $G$ on dual coordinate system $\xi$ is given by
$G_{d}=\frac{1}{\sigma^{4}}\left(\begin{array}[]{cc}\sigma^{2}+2\mu^{2}&-\mu\\ -\mu&\frac{1}{2}\end{array}\right),$
the determinant of this matrix is given by
$DetG_{d}=\frac{1}{2\sigma^{2}}$
In a riemannian manifold, the matrix of information of Fisher in the frame of reference $\xi$ noted $g_{\alpha\beta}$ is given by
$G_{d}^{-1}=\left(\begin{array}[]{cc}\sigma^{2}&2\mu\sigma^{2}\\ 2\mu\sigma^{2}&2\sigma^{4}+4\mu^{2}\sigma^{2}\end{array}\right)$
Proof.
Using theorem 2, we have
$g_{11}(\xi)=\bar{B_{1}}^{1}\bar{B_{1}}^{1}g_{11}(\theta)+\bar{B_{1}}^{2}\bar{B_{1}}^{2}g_{22}(\theta)$
$g_{11}(\xi)=\left(\frac{\partial\theta_{1}}{\partial\xi_{1}}\right)\left(\frac{1}{\sigma^{2}}\right)+\left(\frac{\partial\theta_{2}}{\partial\xi_{1}}\right)\left(\frac{2}{\sigma^{2}}\right)$
because $\theta_{1}=\mu,\;\xi_{1}=\mu$
$g_{11}(\xi)$ $=$ $\frac{1}{\sigma^{2}}+\frac{2\mu^{2}}{\sigma^{4}}$
$=$ $\frac{1}{\sigma^{2}}\left(\frac{\sigma^{2}+2\mu^{2}}{\sigma^{2}}\right)$
$=$ $\frac{1}{\sigma^{4}}\left(\sigma^{2}+2\mu^{2}\right)$
then we have
$g_{22}(\xi)$ $=$ $\bar{B_{2}}^{1}\bar{B_{2}}^{1}g_{11}(\theta)+\bar{B_{2}}^{2}\bar{B_{2}}^{2}g_{22}(\theta)$
$g_{22}(\xi)$ $=$ $\left(\bar{B_{2}}^{1}\right)^{2}g_{11}(\theta)+\left(\bar{B_{2}}^{2}\right)^{2}g_{22}(\theta)$
$g_{22}(\xi)$ $=$ $\left(\bar{B_{2}}^{1}\right)^{2}\left(\frac{1}{\sigma^{2}}\right)+\left(\bar{B_{2}}^{2}\right)^{2}\left(\frac{2}{\sigma^{2}}\right)$
$g_{22}(\xi)$ $=$ $\left(\frac{1}{2\sigma}\right)^{2}\left(\frac{2}{\sigma^{2}}\right)$
$=$ $\frac{2}{\sigma^{4}}$
$i.e.\;g_{22}(\xi)$ $=\frac{1}{2\sigma^{4}}$
$g_{12}(\xi)$ $=$ $\bar{B_{1}}^{1}\bar{B_{2}}^{1}g_{11}(\theta)+\bar{B_{1}}^{2}\bar{B_{2}}^{2}g_{22}(\theta)$
$=$ $\frac{-\mu}{\sigma^{4}}$
$g_{21}(\xi)$ $=$ $\bar{B_{2}}^{1}\bar{B_{1}}^{1}g_{11}(\theta)+\bar{B_{2}}^{1}\bar{B_{1}}^{2}g_{12}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{1}g_{21}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{2}g_{22}(\theta)$
$=$ $\bar{B_{2}}^{1}\bar{B_{1}}^{1}g_{11}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{2}g_{22}(\theta)$
$=$ $\frac{-\mu}{\sigma^{4}}$
We have the result ∎
4 Torsion and curvature tensor in $\xi$-coordinates system.
Theorem 4.
Let $\xi=\left(\mu,\mu^{2}+\sigma^{2}\right)$ the dual coordinate system of $\theta-$ coordinate system on gaussian manifold. The torsion tensor is not free.
Proof.
According to Amari’s [2], we have Christoffel’s coefficient in $\theta$-coordinate system on gaussian statistical manifold
$\Gamma^{(1)}_{111}(\theta)=\Gamma^{(1)}_{112}(\theta)=\Gamma^{(1)}_{221}(\theta)=\Gamma^{(1)}_{122}(\theta)=\Gamma^{(1)}_{212}(\theta)=0$
$\Gamma^{(1)}_{121}(\theta)=\Gamma^{(1)}_{211}(\theta)=-\frac{2}{\sigma^{3}},\;\Gamma^{(1)}_{222}(\theta)=-\frac{6}{\sigma^{3}}.$
Calculation of the coefficients of christoffel’s in the system of coordinates $\xi$.
$\Gamma^{1}_{111}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{1}}^{i}\bar{B_{1}}^{j}\bar{B_{1}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{1}}^{1}\bar{B_{1}}^{1}\bar{B_{1}}^{1}\Gamma_{111}(\theta)+\bar{B_{1}}^{1}\bar{B_{1}}^{1}\bar{B_{1}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{1}}^{1}\bar{B_{1}}^{2}\bar{B_{1}}^{1}\Gamma_{121}(\theta)+\bar{B_{1}}^{1}\bar{B_{1}}^{2}\bar{B_{1}}^{2}\Gamma_{122}(\theta)+$
$+\bar{B_{1}}^{2}\bar{B_{1}}^{1}\bar{B_{1}}^{1}\Gamma_{211}(\theta)+\bar{B_{1}}^{2}\bar{B_{1}}^{1}\bar{B_{1}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{1}}^{2}\bar{B_{1}}^{2}\bar{B_{1}}^{1}\Gamma_{221}(\theta)+\bar{B_{1}}^{2}\bar{B_{1}}^{2}\bar{B_{1}}^{2}\Gamma_{222}(\theta)$
$=$ $-\frac{\mu}{\sigma}.(\frac{-2}{\sigma^{3}})-\frac{\mu}{\sigma}.(\frac{-2}{\sigma^{3}})+\frac{-\mu^{3}}{\sigma^{3}}.(\frac{-6}{\sigma^{3}})$
$=$ $\frac{4\mu\sigma^{2}+6\mu^{3}}{\sigma^{6}}$
$\Gamma_{111}(\xi)$ $=$ $\frac{4\mu\sigma^{2}+6\mu^{3}}{\sigma^{6}}$
By analogy with calculation preceding one a:
$\Gamma_{112}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{1}}^{i}\bar{B_{1}}^{j}\bar{B_{2}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{1}}^{1}\bar{B_{1}}^{1}\bar{B_{2}}^{1}\Gamma_{111}(\theta)+\bar{B_{1}}^{1}\bar{B_{1}}^{1}\bar{B_{2}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{1}}^{1}\bar{B_{1}}^{2}\bar{B_{2}}^{1}\Gamma_{121}(\theta)+\bar{B_{1}}^{1}\bar{B_{1}}^{2}\bar{B_{2}}^{2}\Gamma_{122}(\theta)+$
$\bar{B_{1}}^{2}\bar{B_{1}}^{1}\bar{B_{2}}^{1}\Gamma_{211}(\theta)+\bar{B_{1}}^{2}\bar{B_{1}}^{1}\bar{B_{2}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{1}}^{2}\bar{B_{1}}^{2}\bar{B_{2}}^{1}\Gamma_{221}(\theta)+\bar{B_{1}}^{2}\bar{B_{1}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$=$ $\bar{B_{1}}^{2}\bar{B_{1}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$=$ $\frac{\mu^{2}}{\sigma^{2}}.(\frac{1}{2\sigma})(\frac{-6}{\sigma^{3}})$
$\Gamma_{112}(\xi)$ $=$ $-\frac{3\mu^{2}}{\sigma^{6}}$
In the same way,
$\Gamma_{221}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{2}}^{i}\bar{B_{2}}^{j}\bar{B_{1}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{2}}^{1}\bar{B_{2}}^{1}\bar{B_{1}}^{1}\Gamma_{111}(\theta)+\bar{B_{2}}^{1}\bar{B_{2}}^{1}\bar{B_{1}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{2}}^{1}\bar{B_{2}}^{2}\bar{B_{1}}^{1}\Gamma_{121}(\theta)+\bar{B_{2}}^{1}\bar{B_{2}}^{2}\bar{B_{1}}^{2}\Gamma_{122}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{2}}^{1}\bar{B_{1}}^{1}\Gamma_{211}(\theta)+\bar{B_{2}}^{2}\bar{B_{2}}^{1}\bar{B_{1}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{1}}^{1}\Gamma_{221}(\theta)+\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{1}}^{2}\Gamma_{222}(\theta)$
$=$ $\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{1}}^{2}\Gamma_{222}(\theta)$
$\Gamma_{221}(\xi)$ $=$ $\frac{3\mu}{2\sigma^{6}}$
$\Gamma_{212}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{2}}^{i}\bar{B_{1}}^{j}\bar{B_{2}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{2}}^{1}\bar{B_{1}}^{1}\bar{B_{2}}^{1}\Gamma_{111}(\theta)+\bar{B_{2}}^{1}\bar{B_{1}}^{1}\bar{B_{2}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{2}}^{1}\bar{B_{1}}^{2}\bar{B_{2}}^{1}\Gamma_{121}(\theta)+\bar{B_{2}}^{1}\bar{B_{1}}^{2}\bar{B_{2}}^{2}\Gamma_{122}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{1}}^{1}\bar{B_{2}}^{1}\Gamma_{211}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{1}\bar{B_{2}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{1}}^{2}\bar{B_{2}}^{1}\Gamma_{221}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$\Gamma_{212}(\xi)$ $=$ $\frac{3\mu}{2\sigma^{6}}$
In the same way in a similar way,
$\Gamma_{121}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{1}}^{i}\bar{B_{2}}^{j}\bar{B_{1}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{1}}^{1}\bar{B_{2}}^{1}\bar{B_{1}}^{1}\Gamma_{111}(\theta)+\bar{B_{1}}^{1}\bar{B_{2}}^{1}\bar{B_{1}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{1}}^{1}\bar{B_{2}}^{2}\bar{B_{1}}^{1}\Gamma_{121}(\theta)+\bar{B_{1}}^{1}\bar{B_{2}}^{2}\bar{B_{1}}^{2}\Gamma_{122}(\theta)+$
$\bar{B_{1}}^{2}\bar{B_{2}}^{1}\bar{B_{1}}^{1}\Gamma_{211}(\theta)+\bar{B_{1}}^{2}\bar{B_{2}}^{1}\bar{B_{1}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{1}}^{2}\bar{B_{2}}^{2}\bar{B_{1}}^{1}\Gamma_{221}(\theta)+\bar{B_{1}}^{2}\bar{B_{2}}^{2}\bar{B_{1}}^{2}\Gamma_{222}(\theta)$
$=$ $\bar{B_{1}}^{1}\bar{B_{2}}^{2}\bar{B_{1}}^{1}\Gamma_{121}(\theta)+\bar{B_{1}}^{2}\bar{B_{2}}^{2}\bar{B_{1}}^{2}\Gamma_{222}(\theta)$
$\ \Gamma_{121}(\xi)$ $=$ $-\frac{1}{\sigma^{4}}-\frac{3\mu^{2}}{\sigma^{6}}$
In the same way one obtains
$\Gamma_{211}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{2}}^{i}\bar{B_{1}}^{j}\bar{B_{1}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{2}}^{1}\bar{B_{1}}^{1}\bar{B_{1}}^{1}\Gamma_{111}(\theta)+\bar{B_{2}}^{1}\bar{B_{1}}^{1}\bar{B_{1}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{2}}^{1}\bar{B_{1}}^{2}\bar{B_{1}}^{1}\Gamma_{121}(\theta)+\bar{B_{2}}^{1}\bar{B_{1}}^{2}\bar{B_{1}}^{2}\Gamma_{122}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{1}}^{1}\bar{B_{1}}^{1}\Gamma_{211}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{1}\bar{B_{1}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{1}}^{2}\bar{B_{1}}^{1}\Gamma_{221}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{2}\bar{B_{1}}^{2}\Gamma_{222}(\theta)$
$=$ $\bar{B_{2}}^{2}\bar{B_{1}}^{1}\bar{B_{1}}^{1}\Gamma_{211}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{2}\bar{B_{1}}^{2}\Gamma_{222}(\theta)$
$\Gamma_{211}(\xi)$ $=$ $-\frac{1}{\sigma^{4}}-\frac{3\mu^{2}}{\sigma^{6}}$
$\Gamma_{122}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{1}}^{i}\bar{B_{2}}^{j}\bar{B_{2}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{1}}^{1}\bar{B_{2}}^{1}\bar{B_{2}}^{1}\Gamma_{111}(\theta)+\bar{B_{1}}^{1}\bar{B_{2}}^{1}\bar{B_{2}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{1}}^{1}\bar{B_{2}}^{2}\bar{B_{2}}^{1}\Gamma_{121}(\theta)+\bar{B_{1}}^{1}\bar{B_{2}}^{2}\bar{B_{2}}^{2}\Gamma_{122}(\theta)+$
$\bar{B_{1}}^{2}\bar{B_{2}}^{1}\bar{B_{2}}^{1}\Gamma_{211}(\theta)+\bar{B_{1}}^{2}\bar{B_{2}}^{1}\bar{B_{2}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{1}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{1}\Gamma_{221}(\theta)+\bar{B_{1}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$=$ $\bar{B_{1}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$\Gamma_{122}(\xi)$ $=$ $\frac{3\mu}{2\sigma^{6}}$
$\Gamma_{222}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{2}}^{i}\bar{B_{2}}^{j}\bar{B_{2}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{2}}^{1}\bar{B_{2}}^{1}\bar{B_{2}}^{1}\Gamma_{111}(\theta)+\bar{B_{2}}^{1}\bar{B_{2}}^{1}\bar{B_{2}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{2}}^{1}\bar{B_{2}}^{2}\bar{B_{2}}^{1}\Gamma_{121}(\theta)+\bar{B_{2}}^{1}\bar{B_{2}}^{2}\bar{B_{2}}^{2}\Gamma_{122}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{2}}^{1}\bar{B_{2}}^{1}\Gamma_{211}(\theta)+\bar{B_{2}}^{2}\bar{B_{2}}^{1}\bar{B_{2}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{1}\Gamma_{221}(\theta)+\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$=$ $\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{1}\Gamma_{221}(\theta)+\bar{B_{2}}^{2}\bar{B_{2}}^{1}\bar{B_{2}}^{2}\Gamma_{212}(\theta)+\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$=$ $\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$\Gamma_{222}(\xi)$ $=$ $-\frac{3}{4\sigma^{6}}$
Using 7, we have the result. ∎
Theorem 5.
The curvature tensor in $\xi$- coordinates system of gaussian distribution is given by
$R_{1111}(\xi)$ $=$ $R_{2222}(\xi)=R_{1122}(\xi)=R_{1112}(\xi)=R_{2121}(\xi)=R_{2211}(\xi)=0$
$R_{2221}(\xi)=R_{2212}(\xi)=R_{2122}(\xi)=R_{1121}(\xi)=R_{1211}(\xi)=0$
$R_{1221}(\xi)$ $=$ $\frac{\mu}{\sigma^{6}},\;R_{2112}(\xi)=\frac{6\sigma^{8}+3\mu\sigma^{6}+6\mu^{2}\sigma^{6}+6\mu^{2}}{4\sigma^{14}},\;R_{1212}(\xi)=\frac{1}{2\sigma^{6}}$
$R_{2111}(\xi)$ $=$ $\frac{-2\sigma^{8}+2\mu^{2}\sigma^{6}+46\mu^{2}\sigma^{8}+24\mu^{4}\sigma^{6}+12\mu^{2}\sigma^{2}-4\mu\sigma^{2}+6\sigma^{10}+12\mu^{4}-3\mu^{2}+9\mu\sigma^{8}+18\mu^{3}\sigma^{6}}{2\sigma^{14}}$
and the scalar curve is given by
$\mathcal{K}$ $=$ $\frac{16\mu^{2}\sigma^{8}+12\mu^{4}+6\sigma^{10}-\mu\sigma^{8}-2\mu^{2}\sigma^{2}+10\mu^{3}\sigma^{6}+92\mu^{3}\sigma^{8}+48\mu^{5}\sigma^{6}}{4\sigma^{10}}$
$+\frac{12\mu\sigma^{10}+24\mu^{5}-6\mu^{3}}{4\sigma^{10}}$
Proof.
According to Amari [2], we have
$\Gamma_{\alpha\beta}^{\gamma}(\xi)=\frac{1}{2}g^{\gamma\lambda}(\xi)\left(\frac{\partial g_{\beta\lambda}(\xi)}{\partial\xi_{\alpha}}+\frac{\partial g_{\gamma\alpha}(\xi)}{\partial\xi_{\beta}}-\frac{\partial g_{\alpha\beta}(\xi)}{\partial\xi_{\gamma}}\right)$
with $\alpha,\;\beta,\;\gamma,\;\lambda\in\left\{1,2\right\}$. So, we have the coefficients following
$\Gamma_{11}^{1}(\xi)$ $=$ $\frac{1}{2}g^{11}(\xi)\left(\frac{\partial g_{11}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{11}(\xi)}{\partial\xi_{1}}-\frac{\partial g_{11}(\xi)}{\partial\xi_{1}}\right)+\frac{1}{2}g^{12}(\xi)\frac{\partial g_{12}(\xi)}{\partial\xi_{1}}$
$\Gamma_{11}^{1}(\xi)$ $=$ $\frac{\mu}{\sigma^{2}}.$
$\Gamma_{22}^{2}(\xi)$ $=$ $\frac{1}{2}g^{21}(\xi)\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}+\frac{1}{2}g^{22}(\xi)\frac{\partial g_{22}(\xi)}{\partial\xi_{2}}$
$\Gamma_{22}^{2}(\xi)$ $=$ $0$
$\Gamma_{12}^{2}(\xi)$ $=$ $\frac{1}{2}g^{21}(\xi)\left(\frac{\partial g_{21}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}\right)+\frac{1}{2}g^{22}(\xi)\left(\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}\right)$
$\Gamma_{12}^{2}(\xi)$ $=$ $-\frac{\mu}{\sigma^{2}}.$
$\Gamma_{21}^{2}(\xi)$ $=$ $\frac{1}{2}g^{21}(\xi)\left(\frac{\partial g_{11}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}-\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}\right)+\frac{1}{2}g^{22}(\xi)\left(\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}+\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}-\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}\right)$
$\Gamma_{21}^{2}(\xi)$ $=$ $\frac{\mu}{\sigma^{2}}.$
$\Gamma_{11}^{2}(\xi)$ $=$ $\frac{1}{2}g^{21}(\xi)\left(\frac{\partial g_{11}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{21}(\xi)}{\partial\xi_{1}}-\frac{\partial g_{11}(\xi)}{\partial\xi_{2}}\right)+\frac{1}{2}g^{22}(\xi)\left(\frac{\partial g_{12}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{21}(\xi)}{\partial\xi_{1}}-\frac{\partial g_{11}(\xi)}{\partial\xi_{2}}\right)$
$\Gamma_{11}^{2}(\xi)$ $=$ $\frac{-2\mu^{2}+\mu-3\sigma^{8}-6\mu^{2}\sigma^{6}}{\sigma^{8}}.$
$\Gamma_{21}^{1}(\xi)$ $=$ $\frac{1}{2}g^{11}(\xi)\left(\frac{\partial g_{11}(\xi)}{\partial\xi_{2}}+\frac{\partial g_{12}(\xi)}{\partial\xi_{1}}-\frac{\partial g_{21}(\xi)}{\partial\xi_{1}}\right)+\frac{1}{2}g^{12}(\xi)\left(\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}+\frac{\partial g_{12}}{\partial\xi_{1}}-\frac{\partial g_{21}(\xi)}{\partial\xi_{1}}\right)$
$\Gamma_{21}^{1}(\xi)$ $=$ $\frac{1}{2\sigma^{2}}.$
$\Gamma_{22}^{1}(\xi)$ $=$ $\frac{1}{2}g^{11}(\xi)\left(\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}+\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}\right)+\frac{1}{2}g^{12}(\xi)\left(\frac{\partial g_{22}(\xi)}{\partial\xi_{2}}+\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}\right)$
$\Gamma_{22}^{1}(\xi)$ $=$ $0$
$\Gamma_{12}^{1}(\xi)$ $=$ $\frac{1}{2}g^{11}(\xi)\left(\frac{\partial g_{21}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{11}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{12}(\xi)}{\partial\xi_{1}}\right)+\frac{1}{2}g^{12}(\xi)\left(\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{11}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{1^{\prime}2^{\prime}}}{\partial\xi^{1}}\right)$
$\Gamma_{12}^{1}(\xi)$ $=$ $8\mu^{2}\sigma^{4}+\frac{4\mu+1}{2\sigma^{2}}.$
According to Amari [2], the tensor of Riemanienne curvature of Christoffel is given by
$R_{\alpha\beta\gamma\eta}$ $=$ $<R(\partial_{\xi_{\alpha}},\partial_{\xi_{\beta}},\partial_{\xi_{\gamma}}),\partial_{\xi_{\eta}}>$ (13)
we have
$R(\partial_{\xi_{\alpha}},\partial_{\xi_{\beta}},\partial_{\xi_{\gamma}})$ $=$ $\left(\nabla_{\partial_{\xi_{\alpha}}}\nabla_{\partial_{\xi_{\beta}}}-\nabla_{\partial_{\xi_{\beta}}}\nabla_{\partial_{\xi_{\alpha}}}\right)\partial_{\xi_{\alpha}}-\nabla_{(\partial_{\xi_{\alpha}}\partial_{\xi_{\beta}}-\partial_{\xi_{\beta}}\partial_{\xi_{\alpha}})}\partial_{\xi_{\gamma}}$
$=$ $\left(\nabla_{\partial_{\xi_{\alpha}}}\nabla_{\partial_{\xi_{\beta}}}(\partial_{\xi_{\gamma}})-\nabla_{\partial_{\xi_{\beta}}}\nabla_{\partial_{\xi_{\alpha}}}(\partial_{\xi_{\gamma}})\right)-\nabla_{\partial_{\xi_{\alpha}}\partial_{\xi_{\beta}}}(\partial_{\xi_{\gamma}})+\nabla_{\partial_{\xi_{\beta}}\partial_{\xi_{\alpha}}}(\partial_{\xi_{\gamma}})$
$=$ $\left(\nabla_{\partial_{\xi_{\alpha}}}\Gamma_{\beta\gamma}^{s^{\prime}}\partial_{\xi_{s^{\prime}}}-\nabla_{\partial_{\xi_{\beta}}}\Gamma_{\alpha\gamma}^{s^{\prime}}\partial_{\xi_{s^{\prime}}}\right)-\partial_{\alpha}\nabla_{\partial_{\xi_{\beta}}}(\partial_{\xi_{\gamma}})+\partial_{\xi_{\beta}}\nabla_{\partial_{\xi_{\alpha}}}(\partial_{\xi_{\gamma}})$
$=$ $\left(\partial_{\xi_{\alpha}}\Gamma_{\beta\gamma}^{s^{\prime}}\partial_{\xi_{s^{\prime}}}+\Gamma_{\beta\gamma}^{s^{\prime}}\nabla_{\partial_{\xi_{\alpha}}}\partial_{\xi_{s^{\prime}}}\right)-\left(\partial_{\xi_{\beta}}\Gamma_{\alpha\gamma}^{s^{\prime}}\partial_{\xi_{s^{\prime}}}+\Gamma_{\alpha\gamma}^{s^{\prime}}\nabla_{\partial_{\xi_{\beta}}}\partial_{\xi_{s^{\prime}}}\right)$
$-\partial_{\xi_{\alpha}}\nabla_{\partial_{\xi_{\beta}}}(\partial_{\xi_{\gamma}})+\partial_{\xi_{\beta}}\nabla_{\partial_{\xi_{\alpha}}}(\partial_{\xi_{\gamma}})$
$=$ $\left(\partial_{\xi_{\alpha}}\Gamma_{\beta\gamma}^{s^{\prime}}\partial_{\xi_{s^{\prime}}}+\Gamma_{\beta\gamma}^{s^{\prime}}\nabla_{\partial_{\xi_{\alpha}}}\partial_{\xi_{s^{\prime}}}\right)-\left(\partial_{\xi_{\beta}}\Gamma_{\alpha\gamma}^{s^{\prime}}\partial_{\xi_{s^{\prime}}}+\Gamma_{\alpha\gamma}^{s^{\prime}}\nabla_{\partial_{\xi_{\beta}}}\partial_{\xi_{s^{\prime}}}\right)$
$-\partial_{\xi_{\alpha}}\Gamma_{\beta\gamma}^{r^{\prime}}\partial_{\xi_{r^{\prime}}}+\partial_{\xi_{\beta}}\Gamma_{\alpha\gamma}^{r^{\prime}}\partial_{\xi_{r^{\prime}}}$
While replacing in (13) we obtain
$R_{\alpha\beta\gamma\eta}$ $=$ $\left(\partial_{\xi_{\alpha}}\Gamma^{s^{\prime}}_{\beta\gamma}-\partial_{\xi_{\beta}}\Gamma^{s^{\prime}}_{\alpha\gamma}\right)g_{s^{\prime}\eta}-\left(\Gamma^{r^{\prime}}_{\beta\gamma}\Gamma_{\alpha r^{\prime}\eta}-\Gamma^{r^{\prime}}_{\alpha\gamma}\Gamma_{\beta r^{\prime}\eta}\right)$
$+<\Gamma^{s^{\prime}}_{\beta\gamma}\nabla_{\partial_{\xi_{\alpha}}}\partial_{\xi_{s^{\prime}}},\partial_{\xi_{\eta}}>-<\Gamma^{s^{\prime}}_{\alpha\gamma}\nabla_{\partial_{\xi_{\beta}}}\partial_{\xi_{s^{\prime}}},\partial_{\xi_{\eta}}>$
with $\alpha,\beta,\gamma,\eta,s^{\prime}\in\left\{1,2\right\}$ so,
$R_{1111}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{11}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{11}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{111}(\xi)\Gamma_{11}^{1}(\xi)-\Gamma_{111}(\xi)\Gamma_{11}^{1}(\xi)\right)-\left(\Gamma_{121}(\xi)\Gamma_{11}^{2}(\xi)-\Gamma_{121}(\xi)\Gamma_{11}^{2}(\xi)\right)$
$+<\Gamma_{11}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{1}}>+<\Gamma_{11}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{1}}>$
$-<\Gamma_{11}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{1}}>-<\Gamma_{11}^{2}\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{1}}),\partial_{\xi_{1}}>$
$=$ 00 0
$R_{2222}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{22}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{22}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{22}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{22}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{222}(\xi)\Gamma_{22}^{1}(\xi)-\Gamma_{222}(\xi)\Gamma_{22}^{1}(\xi)\right)-\left(\Gamma_{222}(\xi)\Gamma_{22}^{2}(\xi)-\Gamma_{222}(\xi)\Gamma_{22}^{2}(\xi)\right)$
$+<\Gamma_{22}^{1}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>+<\Gamma_{22}^{2}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$-<\Gamma_{22}^{1}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>-<\Gamma_{22}^{2}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$=$ 00 0
$R_{1122}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{12}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{12}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{12}^{2}(\xi)-\partial_{1^{\prime}}\Gamma_{12}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{112}(\xi)\Gamma_{12}^{1}(\xi)-\Gamma_{112}(\theta)\Gamma_{12}^{1}(\xi)\right)-\left(\Gamma_{122}(\xi)\Gamma_{12}^{2}(\xi)-\Gamma_{122}(\xi)\Gamma_{12}^{2}(\xi)\right)$
$+<\Gamma_{12}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>+<\Gamma_{12}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$-<\Gamma_{12}^{1}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>-<\Gamma_{12}^{2}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$=$ $0$
$R_{1112}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{11}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{11}^{2}(\xi)-\partial_{1^{\prime}}\Gamma_{11}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{112}(\xi)\Gamma_{11}^{1}(\xi)-\Gamma_{112}(\xi)\Gamma_{11}^{1}(\xi)\right)-\left(\Gamma_{122}(\xi)\Gamma_{11}^{2}(\xi)-\Gamma_{122}(\xi)\Gamma_{11}^{2}(\xi)\right)$
$+<\Gamma_{11}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>+<\Gamma_{11}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$-<\Gamma_{11}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>-<\Gamma_{11}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$=$ 00 0
$R_{2121}(\xi)$ $=$ $\left(\partial_{2^{\prime}}\Gamma_{12}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{22}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{12}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{22}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{211}(\xi)\Gamma_{12}^{1}(\xi)-\Gamma_{111}(\xi)\Gamma_{22}^{1}(\xi)\right)-\left(\Gamma_{221}(\xi)\Gamma_{12}^{2}(\xi)-\Gamma_{121}(\xi)\Gamma_{22}^{2}(\xi)\right)$
$+<\Gamma_{12}^{1}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{1}}),\partial_{\xi_{1}}>+<\Gamma_{12}^{2}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{2}}),\partial_{\xi_{1}}>$
$-<\Gamma_{22}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{1}}>-<\Gamma_{22}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{1}}>$
$=$ 00 0
$R_{2211}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{21}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{21}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{21}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{21}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{211}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{211}\Gamma_{21}^{1}(\xi)\right)-\left(\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)\right)$
$+\Gamma_{211}(\xi)\Gamma_{21}^{1}(\xi)+\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{211}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)$
$=$ $0$
$R_{2221}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{22}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{22}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{22}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{22}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{211}(\xi)\Gamma_{22}^{1}(\xi)-\Gamma_{211}(\xi)\Gamma_{22}^{1}(\xi)\right)-\left(\Gamma_{211}(\xi)\Gamma_{22}^{1}(\xi)-\Gamma_{211}(\xi)\Gamma_{22}^{1}(\xi)\right)$
$+\Gamma_{211}(\xi)\Gamma_{22}^{1}(\xi)+\Gamma_{221}\Gamma_{22}^{2}(\xi)-\Gamma_{211}(\xi)\Gamma_{22}^{1}(\xi)-\Gamma_{221}(\xi)\Gamma_{22}^{2}(\xi)$
$=$ $0$
$R_{2212}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{21}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{21}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{21}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{21}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{212}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{212}(\xi)\Gamma_{21}^{1}(\xi)\right)-\left(\Gamma_{222}\Gamma_{21}^{2}(\xi)-\Gamma_{222}(\xi)\Gamma_{21}^{2}(\xi)\right)$
$+\Gamma_{211}\Gamma_{21}^{1}(\xi)+\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{211}\Gamma_{21}^{1}(\xi)-\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)$
$=$ $0$
$R_{2122}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{12}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{22}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{12}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{22}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{212}(\xi)\Gamma_{12}^{1}(\xi)-\Gamma_{112}(\xi)\Gamma_{22}^{1}(\xi)\right)-\left(\Gamma_{222}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{122}(\xi)\Gamma_{22}^{2}(\xi)\right)$
$+\Gamma_{212}(\xi)\Gamma_{12}^{1}(\xi)+\Gamma_{222}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{211}\Gamma_{22}^{1}(\xi)-\Gamma_{122}(\xi)\Gamma_{22}^{2}(\xi)$
$=$ $0$
$R_{1221}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{22}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{12}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{22}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{12}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{111}\Gamma_{22}^{1}(\xi)-\Gamma_{211}(\xi)\Gamma_{21}^{1}(\xi)\right)-\left(\Gamma_{121}(\xi)\Gamma_{22}^{2}(\xi)-\Gamma_{221}(\xi)\Gamma_{12}^{2}(\xi)\right)$
$+\Gamma_{111}(\xi)\Gamma_{21}^{1}(\xi)+\Gamma_{121}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{211}(\xi)\Gamma_{12}^{1}(\xi)-\Gamma_{221}\Gamma_{12}^{2}(\xi)$
$=$ $\Gamma_{111}(\xi)\Gamma_{21}^{1}(\xi)+\Gamma_{121}(\xi)\Gamma_{21}^{2}(\xi)$
$=$ $\frac{\mu}{\sigma^{6}}.$
$R_{2112}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{11}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{21}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{11}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{21}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{212}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{112}(\xi)\Gamma_{21}^{1}(\xi)\right)-\left(\Gamma_{222}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{222}(\xi)\Gamma_{21}^{2}(\xi)\right)$
$+\Gamma_{212}(\xi)\Gamma_{11}^{1}(\xi)+\Gamma_{222}(\xi)\Gamma_{11}^{2}(\xi)-\Gamma_{112}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{122}(\xi)\Gamma_{21}^{2}(\xi)$
$=$ $\frac{6\sigma^{8}+3\mu\sigma^{6}+6\mu^{2}\sigma^{6}+6\mu^{2}}{4\sigma^{14}}.$
$R_{1212}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{21}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{11}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{12}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{11}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{112}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{212}(\xi)\Gamma_{11}^{1}(\xi)\right)-\left(\Gamma_{122}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{222}(\xi)\Gamma_{11}^{2}(\xi)\right)$
$+<\Gamma_{21}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>+<\Gamma_{21}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$-<\Gamma_{11}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>-<\Gamma_{11}^{2}\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$=\left(\partial_{\xi_{1}}\Gamma_{21}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{11}^{1}\right)g_{12}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{12}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{11}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{112}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{212}\Gamma_{11}^{1}\right)-\left(\Gamma_{122}\Gamma_{21}^{2}-\Gamma_{222}(\xi)\Gamma_{11}^{2}(\xi)\right)$
$+\Gamma_{112}(\xi)\Gamma_{21}^{1}(\xi)+\Gamma_{122}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{212}(\xi)\Gamma_{11}^{1}(\xi)-\Gamma_{222}(\xi)\Gamma_{11}^{2}(\xi)$
$=$ $\frac{1}{2\sigma^{6}}.$
$R_{1211}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{21}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{11}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{21}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{11}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{111}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{211}(\xi)\Gamma_{11}^{1}(\xi)\right)-\left(\Gamma_{121}\Gamma_{21}^{2}(\xi)-\Gamma_{221}(\xi)\Gamma_{11}^{2}(\xi)\right)$
$+\Gamma_{111}(\xi)\Gamma_{21}^{1}(\xi)+\Gamma_{121}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{221}(\xi)\Gamma_{11}^{2}(\xi)-\Gamma_{211}(\xi)\Gamma_{11}^{1}(\xi)$
$=$ $0$
$R_{1121}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{12}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{12}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{12}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{12}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{111}(\xi)\Gamma_{12}^{1}(\xi)-\Gamma_{111}(\xi)\Gamma_{12}^{1}(\xi)\right)-\left(\Gamma_{121}\Gamma_{12}^{2}(\xi)-\Gamma_{121}(\xi)\Gamma_{12}^{2}(\xi)\right)$
$+\Gamma_{111}(\xi)\Gamma_{12}^{1}(\xi)+\Gamma_{121}(\xi)\Gamma_{12}^{2}(\xi)-\Gamma_{111}(\xi)\Gamma_{12}^{1}(\xi)-\Gamma_{121}(\xi)\Gamma_{12}^{2}(\xi)$
$=$ $0$
$R_{2111}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{11}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{11}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{211}(\xi)\Gamma_{11}^{1}(\xi)-\Gamma_{112}(\xi)\Gamma_{21}^{1}(\xi)\right)-\left(\Gamma_{221}(\xi)\Gamma_{11}^{2}(\xi)-\Gamma_{122}(\xi)\Gamma_{21}^{2}(\xi)\right)$
$+<\Gamma_{11}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{1}}>+<\Gamma_{11}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{1}}>$
$-<\Gamma_{21}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>-<\Gamma_{21}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$=$ $\left(\partial_{2^{\prime}}\Gamma_{11}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{11}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{211}\Gamma_{11}^{1}(\xi)-\Gamma_{112}\Gamma_{21}^{1}(\xi)\right)-\left(\Gamma_{221}(\xi)\Gamma_{11}^{2}(\xi)-\Gamma_{122}(\xi)\Gamma_{21}^{2}(\xi)\right)$
$+\Gamma_{11}^{1}(\xi)\Gamma_{111}(\xi)+\Gamma_{11}^{1}(\xi)\Gamma_{121}(\xi)-\Gamma_{21}^{1}(\xi)-\Gamma_{21}^{2}(\xi)\Gamma_{122}(\xi)$
$=$ $\frac{-2\sigma^{8}+2\mu^{2}\sigma^{6}+46\mu^{2}\sigma^{8}+24\mu^{4}\sigma^{6}+12\mu^{2}\sigma^{2}-4\mu\sigma^{2}+6\sigma^{10}+12\mu^{4}-3\mu^{2}+9\mu\sigma^{8}+18\mu^{3}\sigma^{6}}{2\sigma^{14}}$
Using the definition 11, we have
$\mathcal{K}$ $=$ $\frac{1}{2}R_{1111}(\xi)g^{11}(\xi)g^{11}(\xi)+\frac{1}{2}R_{1112}(\xi)g^{12}(\xi)g^{11}(\xi)+\frac{1}{2}R_{1121}(\xi)g^{11}(\xi)g^{12}(\xi)$
$+\frac{1}{2}R_{1122}(\xi)g^{12}(\xi)g^{12}(\xi)+\frac{1}{2}R_{1211}(\xi)g^{11}(\xi)g^{21}(\xi)+\frac{1}{2}R_{1212}(\xi)g^{12}(\xi)g^{21}(\xi)$
$+\frac{1}{2}R_{1221}(\xi)g^{11}(\xi)g^{22}(\xi)+\frac{1}{2}R_{1222}(\xi)g^{12}(\xi)g^{22}(\xi)+\frac{1}{2}R_{2111}(\xi)g^{21}(\xi)g^{11}(\xi)$
$+\frac{1}{2}R_{2112}(\xi)g^{22}(\xi)g^{11}(\xi)+\frac{1}{2}R_{2121}(\xi)g^{21}(\xi)g^{12}(\xi)+\frac{1}{2}R_{2122}(\xi)g^{22}(\xi)g^{12}(\xi)$
$+\frac{1}{2}R_{2211}(\xi)g^{21}(\xi)g^{21}(\xi)+\frac{1}{2}R_{2212}(\xi)g^{22}g^{21}(\xi)+\frac{1}{2}R_{2221}(\xi)g^{21}(\xi)g^{22}(\xi)$
$+\frac{1}{2}R_{2222}(\xi)g^{22}(\xi)g^{22}(\xi)$
$=$ $\frac{1}{2}R_{1212}(\xi)g^{12}(\xi)g^{21}(\xi)+\frac{1}{2}R_{2111}(\xi)g^{21}(\xi)g^{11}(\xi)+\frac{1}{2}R_{2112}(\xi)g^{22}(\xi)g^{11}(\xi)$
$=$ $\frac{16\mu^{2}\sigma^{12}+12\mu^{4}\sigma^{4}+6\sigma^{14}-\mu\sigma^{12}-2\mu^{2}\sigma^{6}+10\mu^{3}\sigma^{10}+92\mu^{3}\sigma^{12}+48\mu^{5}\sigma^{10}}{4\sigma^{14}}$
$+\frac{12\mu\sigma^{14}+24\mu^{5}\sigma^{4}-6\mu^{3}\sigma^{4}}{4\sigma^{14}}$
Thus the curve will be given by
$\mathcal{K}$ $=$ $\frac{16\mu^{2}\sigma^{8}+12\mu^{4}+6\sigma^{10}-\mu\sigma^{8}-2\mu^{2}\sigma^{2}+10\mu^{3}\sigma^{6}+92\mu^{3}\sigma^{8}+48\mu^{5}\sigma^{6}}{4\sigma^{10}}$
$+\frac{12\mu\sigma^{10}+24\mu^{5}-6\mu^{3}}{4\sigma^{10}}$
5 General conclusion
In conclusion, on a Gaussian manifold precisely in the dual coordinate system the torsion and scalar curvature is non-zero. The notion of duality does not preserve the non-linear properties which are curvature and torsion in the $\theta-$coordinate system.
[…]
Supplementary information This manuscript has no additional data.
[…]
Acknowledgments We would like to thank all the active members of the algebra and geometry research group at the University of Maroua in Cameroon.
Declarations
This article has no conflict of interest to the journal. No financing with a third party.
  • No Funding
  • No Conflict of interest/Competing interests (check journal-specific guidelines for which heading to use)
  • Ethics approval
  • Consent to participate
  • Consent for publication
  • Availability of data and materials
  • Code availability
  • Authors’ contributions

### v2

@@
 Abstract
-In this paper, we study the geometric nonlinearity properties, such as curvature and torsion, in a dual coordinate system of the Riemannian manifold defined by the Gaussian distribution.
-We also give the Amari formulas explicitly in this new coordinate system, which allows us to characterize existing geometric invariants, such as the dual potential function and the Fisher metric.
+This paper explore the differential geometric structure of statistical manifolds defined by Gaussian distributions, focusing on their representation in a dual coordinate system.
+Using Amari’s information geometry framework, we derive explicit expressions for the Fisher information metric and compute the associated torsion and curvature tensors in the dual setting.
+Our results show that, unlike the standard coordinate system where curvature and torsion vanish, the dual system exhibits non zero values for these three geometric quantities.
+This highlights how dual parameterizations can affect the intrinsic geometry of statistical models.
 keywords:
@@
 1 Introduction
-Information geometry, although a very well-known and rich branch of mathematics, has very important tools for studying certain geometric properties such as curvature and torsion, and for studying probability and information through differential geometry for example see [1].
-The main idea is to study nonlinear geometric properties, such as curvature and torsion, in a dual coordinate system on a statistical manifold defined by a Gaussian distribution.
-Several researchers have studied this field, including S.
-I.
-Amari [2], one of the pioneers of information geometry.
-M.
-N.
-Boyom [3], who studies umerical properties of Koszul connections’.
-Koszul and Jean-Louis [4], who study deformations of locally flat connections.
-In [5], the geometric properties of beta distributions are studied.
-Calculating curvature and torsion is important because it allows a statistical manifold to be described globally.
-Curvature is used in fields such as physics.
-Since this manifold is torsion-free and curvature-free in the local coordinate system, can these properties be preserved in the dual coordinate system to conserve the obtained values?
-Research in this area dates back to Rao, who first recognized the importance of the differential geometry approach in statistics.
-He defined the Riemannian metric in terms of the Fisher information matrix, $g^{ij}$, and calculated the geodesic distances between two distributions for various statistical models.
-Amari [2], studies the geometric properties of the Riemannian manifold of a statistical model.
-He has shown that the curvature is entirely contained within the plane, and that the curvature in the $\theta$-coordinate system is constant.
-Building on Amari’s work, Ovidiu Calin[6] also presents the properties of the Gaussian model as a Riemannian manifold and its characteristics in the $\theta$-coordinate system.
-In [7], it then introduces the Riemann curvature tensor, and quickly moves on to sub-manifold theory in order to give the curvature tensor a concrete quantitative interpretation.
-In [8], it is shown that the statistical manifold define by lognormal distribution is flat with respect to a pair of torsion-free dual affine connections $\nabla$ and $\nabla^{*}$.
-Recently it has been observed that the space of fields $T^{l}(\theta)$ is seen as a subspace of the space of second degree trinomial with coefficients in $\sigma^{-3}C^{\infty}(S)$ where $\sigma^{-3}C^{\infty}(S)$ is a modulus $\mathbb{R}$.
-where $\sigma^{-3}C^{\infty}(S)$ is a modulus $\mathbb{R}$.
-In the same order as Amari and Bovai Lobaschevki on the manifold of models we have proved that the curvature in the $\xi$-coordinate system is equal to
+Information geometry is a powerful mathematical framework that applies the methods of differential geometry to the study of statistical models and probability distributions.
+By treating families of probability distributions as points on a manifold, this field provides a geometric interpretation of key statistical concepts such as divergence, estimation, and entropy.
+Among its central tools are the Fisher information metric and dual affine connections, which allow one to study geometric invariants like curvature and torsion within a statistical setting.
+The foundational work of C.R.
+Rao [2], first demonstrated the relevance of differential geometry in statistics by introducing a Riemannian metric defined through the Fisher information matrix, enabling the computation of geodesic distances between distributions.
+Building on this insight, S.
+Amari and H.
+Nagaoka [2], developed a comprehensive theory of information geometry, formalizing the notions of dual coordinate systems and affine connections on statistical manifolds.
+Their theory shows that many statistical models, including the Gaussian family, exhibit flatness and torsion free properties in their natural (exponential or mixture) coordinate systems.
+However, the behavior of geometric properties under dual coordinate transformations remains a subtle and open question.
+While it is well known that the Gaussian manifold is flat and torsion-free in its natural coordinate system $\theta$, it is not always clear whether such nonlinear geometric invariants are preserved when one moves to a dual coordinate system, such as $\xi=(\mu,\mu^{2}+\sigma^{2})$?
+This leads to the central problematic of this work: To what extent does the dual coordinate transformation affect the intrinsic nonlinear geometry (curvature and torsion) of a Gaussian statistical manifold?
+Are these geometric invariants preserved, distorted, or enriched in the dual setting?
+Several researchers have addressed related questions.
+M.N.
+Boyom [1, 3] has explored the geometric and numerical properties of Koszul connections in statistical settings, while J.-L.
+Koszul, and collaborators [4], have studied deformations of locally flat connections.
+In more recent works, in [5], investigated the geometric structure of beta distributions within the framework of information geometry.
+Ovidiu Calin [6], further developed the differential geometric modeling of the Gaussian distribution, providing explicit expressions for curvature and connection coefficients in natural coordinates.
+In addition, J.M.
+Lee [7], offered a comprehensive exposition of Riemannian manifolds and curvature tensors that underpins the theoretical foundation of our investigation.
+It is also worth noting that recent developments have shown the space of vector fields $T^{l}(\theta)$ on a Gaussian manifold can be identified with a submodule of second-degree polynomials with smooth coefficients scaled by $\sigma^{-3}$, i.e., $\sigma^{-3}C^{\infty}(S)[x]^{2}$, offering new algebraic insights into the tangent bundle structure.
+In this work, we extend these investigations by focusing on the Riemannian and affine geometric structures of the Gaussian statistical manifold when expressed in the dual coordinate system $\xi=(\mu,\mu^{2}+\sigma^{2})$.
+We derive explicit formulas for the Fisher information metric, the Christoffel symbols, the torsion tensor, and the Riemann curvature tensor in this dual system.
+We prove that the curvature in the $\xi$-coordinate system is equal to
 $\mathcal{K}$ $=$ $\frac{16\mu^{2}\sigma^{8}+12\mu^{4}+6\sigma^{10}-\mu\sigma^{8}-2\mu^{2}\sigma^{2}+10\mu^{3}\sigma^{6}+92\mu^{3}\sigma^{8}+48\mu^{5}\sigma^{6}}{4\sigma^{10}}$
 $+\frac{12\mu\sigma^{10}+24\mu^{5}-6\mu^{3}}{4\sigma^{10}}$
-we will study the non-linear properties of the manifold in system of the $\xi$- coordinates which contains a gaussian model.
-An extension of this work would then be to formalize to what degree certain geometric information does not vary for the normal law whatever the reference field.
-After the introduction, the first section recall the preliminaries motion on theory of statistical manifold, in section $2$ we determine the Riemannian structure in $\xi$-coordinates system, in section $3$, we determine torsion tensor and curvature tensor.
-At the end, we present the conclusion.
+A key outcome of our analysis is the emergence of a non zero scalar curvature, which contrasts with the flatness observed in the natural coordinates.
+This result suggests that dual parametrization may introduce richer nonlinear geometric features, with implications for the stability and global structure of statistical models.
+The structure of the paper is as follows: Section 2 reviews the basic theoretical framework of statistical manifolds and introduces the Gaussian model.
+Section 3 presents the derivation of the Riemannian metric in the dual $\xi$-coordinate system.
+Section 4 is devoted to computing the torsion and curvature tensors in this setting.
+We conclude with a discussion of the implications of our findings and potential directions for future work, including generalizations to other distributions and geometric invariants.
 2 Preliminaries
-Soit $S=\left\{p_{\theta}(x),\left.\begin{array}[]{ll}\theta\in\Theta&\hbox{}\\ x\in\mathcal{X}&\hbox{}\end{array}\right.\right\}$ be the set of probabilities $p_{\theta}$, parameterized by $\Theta$, open subset of $\mathbb{R}^{n}$; on the sample space $\mathcal{X}\subseteq\mathbb{R}$, with $\theta=\left(\theta_{1},\theta_{2},\dots,\theta_{n}\right)$.
-In [8], the Gaussian model is a family of probabilistic distributions.
-consisting of density functions that are defined as follows $S=\left\{p_{\theta}(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}},\left.\begin{array}[]{ll}\theta\in\mathbb{R}\times\mathbb{R}^{*}_{+}&\hbox{}\\ x\in\mathbb{R}^{*}_{+}&\hbox{}\end{array}\right.\right\}$ with $p_{\theta}(x)$ the gaussian distribution parameterized by $\theta=(\mu,\sigma)$.
+[…]
 2.1 $\theta$-coordinates System and $\xi$-coordinates system..
@@
 2.2 Determination of the characteristic elements of the $\xi$- coordinates.
-Let $S=\left\{p_{\theta}(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}},\left.\begin{array}[]{ll}\theta\in\mathbb{R}\times\mathbb{R}^{*}_{+}&\hbox{}\\ x\in\mathbb{R}^{*}_{+}&\hbox{}\end{array}\right.\right\}$ the gaussian statistical model.
-The tangent space of $l-$representation is given by
-$T^{l}(\theta)$ $=$ =$=\left\{A(x);A(x)=A^{1}\frac{\partial_{1}l(x,\theta)}{\partial\mu}+A^{2}\frac{\partial_{2}l(x,\theta)}{\partial\sigma}\right\}$
+[…]
+$T^{l}(\theta)$ $=$ $=\left\{A(x);A(x)=A^{1}\frac{\partial_{1}l(x,\theta)}{\partial\mu}+A^{2}\frac{\partial_{2}l(x,\theta)}{\partial\sigma}\right\}$
 and using (2.1) we have
@@
 According to [2], it is shown that
-$\mathbb{E}[\partial_{i}l(x,\theta)]$ $=$ $0$ (2)
+$\mathbb{E}[\partial_{i}l(x,\theta)]$ $=$ $0.$ (2)
 Furthermore, it follows from that
@@
 using (2) we obtain
-$\mathbb{E}[A(x)]$ $=$ 00 0
+$\mathbb{E}[A(x)]$ $=$ $0$
 […]
@@
 Denote by
-(¯$(\overline{\mathcal{B}}^{i}_{\alpha})=\left(\begin{array}[]{cc}\frac{\partial\theta_{1}}{\partial\xi_{1}}&\frac{\partial\theta_{1}}{\partial\xi_{2}}\\ \frac{\partial\theta_{2}}{\partial\xi_{1}}&\frac{\partial\theta_{2}}{\partial\xi_{2}}\\ \end{array}\right)=J_{M}^{-1}$
+$(\overline{\mathcal{B}}^{i}_{\alpha})=\left(\begin{array}[]{cc}\frac{\partial\theta_{1}}{\partial\xi_{1}}&\frac{\partial\theta_{1}}{\partial\xi_{2}}\\ \frac{\partial\theta_{2}}{\partial\xi_{1}}&\frac{\partial\theta_{2}}{\partial\xi_{2}}\\ \end{array}\right)=J_{M}^{-1}$
 the inverse of the Jacobian matrix $J_{M}$, and that
@@
 Its $l$-representation is $\partial_{\xi}l(x,\theta)\}=\left\{\partial_{1}l(x,\theta),-\frac{\mu}{\sigma}\partial_{1}l(x,\theta)+\frac{1}{2\sigma}\partial_{2}l(x,\theta)\right\}$
-We have $\left\{\begin{array}[]{ll}\partial_{\xi_{1}}l(x,\theta)=\frac{x-\mu}{\sigma^{2}}&\hbox{}\\ \partial_{\xi_{2}}l(x,\theta)=\frac{(x-\mu)^{2}}{2\sigma^{4}}-\frac{\mu(x-\mu)}{\sigma^{3}}-\frac{1}{2\sigma^{2}}&\hbox{}\end{array}\right.$
+[…]
 Remark 3.
@@
 $T:\mathfrak{X}(S)\times\mathfrak{X}(S)$ $\longrightarrow$ $\mathfrak{X}(S):(A,B)\longmapsto T(A,B)=\nabla_{A}(B)-\nabla_{B}(A)-[A,B]$
-In the case of statistical manifold $S=\left\{p_{\theta}(x),\left.\begin{array}[]{ll}\theta\in\Theta&\hbox{}\\ x\in\mathcal{X}&\hbox{}\end{array}\right.\right\}$. Amari [2], show that
-$\mathbb{E}[\partial_{i}l(x,\theta)]$ $=$ $0$ (5)
-Where $l(x,\theta)$ is the $l-$representation of $T_{\theta}$.
-In this case by taking into consideration the Amari [2], the result affine connection of $T_{\theta}$ on statistical manifold $S=\left\{p_{\theta}(x),\left.\begin{array}[]{ll}\theta\in\Theta&\hbox{}\\ x\in\mathcal{X}&\hbox{}\end{array}\right.\right\}$ is define by
+[…]
+$\mathbb{E}[\partial_{i}l(x,\theta)]$ $=$ $0.$ (5)
+[…]
 $\Gamma_{ijk}:=\mathbb{E}\left[\partial_{i}\partial_{j}l(x,\theta).\partial_{k}l(x,\theta)\right].$ (6)
@@
 The curvature tensor of affine connection $\nabla$ on a $n$-manifold $S$ is a $\mathbb{R}$-trilinear mapping $R:\mathfrak{X}(S)\times\mathfrak{X}(S)\times\mathfrak{X}(S)\longrightarrow\mathfrak{X}(S):(X,Y,Z)\longmapsto R(X,Y)Z$
-such that $R(X,Y)Z=[\nabla_{X},\nabla_{Y}]Z-\nabla_{[X,Y]}Z,$ with $[\nabla_{X},\nabla_{Y}]Z=\nabla_{X}\nabla_{Y}Z-\nabla_{Y}\nabla_{X}Z.$ In a locale coordinates $\theta=\left(\theta_{1},\dots,\theta_{n}\right)$ of $S$, we have $R(\partial_{i},\partial_{j}).\partial_{k}=\sum_{m=1}^{n}R_{ijk}^{m}\partial_{m}$
-where $R_{ijk}^{m}=\frac{\partial\Gamma_{kl}^{m}}{\partial_{j}}-\frac{\partial\Gamma_{jl}^{m}}{\partial_{k}}+\Gamma_{js}^{m}\Gamma_{kl}^{s}-\Gamma_{ks}^{m}\Gamma_{jl}^{s}.$ Taking in consideration the $l-$representation of $T_{\theta}$ on statistical manifold $S=\left\{p_{\theta}(x),\left.\begin{array}[]{ll}\theta\in\Theta&\hbox{}\\ x\in\mathcal{X}&\hbox{}\end{array}\right.\right\}$.
-The family
+such that $R(X,Y)Z=[\nabla_{X},\nabla_{Y}]Z-\nabla_{[X,Y]}Z,$ with $\nabla_{X},\nabla_{Y}]Z=\nabla_{X}\nabla_{Y}Z-\nabla_{Y}\nabla_{X}Z.$ In a locale coordinates $\theta=\left(\theta_{1},\dots,\theta_{n}\right)$ of $S$, we have $R(\partial_{i},\partial_{j}).\partial_{k}=\sum_{m=1}^{n}R_{ijk}^{m}\partial_{m}$
+[…]
 $R_{ijkm}:=\langle R(\partial_{i},\partial_{j},\partial_{k}),\partial_{m}\rangle$ (8)
@@
 $=$ $-\frac{2}{\sigma}\mathbb{E}\left[\partial_{1}l(x,\theta)\right]$
-$=$ 00 0
+$=$ $0$
 $g_{21}(\theta)$ $=$ $-\mathbb{E}\left[\partial_{2}\partial_{1}l(x,\theta)\right]$
@@
 $=$ $\frac{2}{\sigma}\mathbb{E}\left[\partial_{1}l(x,\theta)\right]$
-$=$ 00 0
+$=$ $0$
 So, the metric one associated for a manifold of gaussian model $S$ is given by
@@
 Theorem 2.
-Let $\xi=\left(\mu,\mu^{2}+\sigma^{2}\right)$ the $\xi$-coordinate system on gaussian manifold.
-Let $\left(g_{ij}\right)_{1\leq i,j\leq 2}$ the Riemannian of metric in $\theta-$ coordinate on $S$ given by equation (12), and (¯$(\overline{\mathcal{B}}^{i}_{\alpha})=\left(\begin{array}[]{cc}1&0\\ \frac{-\mu}{\sigma}&\frac{1}{2\sigma}\\ \end{array}\right)$ The inverse of the jacobian matrix.
-The Riemannian of metric $g_{\alpha\beta}(\xi)$, in the $\xi$- co-ordinates is given by
-$g_{\alpha\beta}(\xi)=\bar{B_{\alpha}}^{1}\bar{B_{\beta}}^{1}g_{11}(\theta)+\bar{B_{\alpha}}^{2}\bar{B_{\beta}}^{2}g_{22}(\theta)$
+Let $\xi=\left(\mu,\mu^{2}+\sigma^{2}\right)$ be the dual coordinate system associated with the Gaussian statistical manifold.
+Let $\left(g_{ij}\right)_{1\leq i,j\leq 2}$ denote the Riemannian metric expressed in the natural coordinate system $\theta=(\mu,\sigma)$, as defined by equation (12).
+Let $(\overline{\mathcal{B}}^{i}_{\alpha})$ denote the inverse of the Jacobian matrix of the coordinate transformation from $\theta$ to $\xi$, given explicitly by: $(\overline{\mathcal{B}}^{i}_{\alpha})=\left(\begin{array}[]{cc}1&0\\ \frac{-\mu}{\sigma}&\frac{1}{2\sigma}\\ \end{array}\right)$ Then, the components of the Riemannian metric $g_{\alpha\beta}(\xi)$ in the $\xi$-coordinate system are given by:
+$g_{\alpha\beta}(\xi)=\frac{1}{\sigma^{2}}\bar{B_{\alpha}}^{1}\bar{B_{\beta}}^{1}+\frac{2}{\sigma^{2}}\bar{B_{\alpha}}^{2}\bar{B_{\beta}}^{2}$
 $\alpha,\;\beta\in\left\{1,2\right\}$.
 Proof.
+Let $\theta=(\mu,\sigma)$ denote the natural coordinate system and $\xi=(\mu,\mu^{2}+\sigma^{2})$ the dual coordinate system on the Gaussian manifold.
+From the coordinate transformation, the Jacobian matrix of the map $\theta\mapsto\xi$ is given by:
+$J=\begin{pmatrix}\frac{\partial\xi_{1}}{\partial\mu}&\frac{\partial\xi_{1}}{\partial\sigma}\\ \frac{\partial\xi_{2}}{\partial\mu}&\frac{\partial\xi_{2}}{\partial\sigma}\end{pmatrix}=\begin{pmatrix}1&0\\ 2\mu&2\sigma\end{pmatrix}$,
+whose inverse is:
+$J^{-1}=(\overline{\mathcal{B}}^{i}_{\alpha})=\begin{pmatrix}1&0\\ -\frac{\mu}{\sigma}&\frac{1}{2\sigma}\end{pmatrix}$.
+Let $\{\partial_{1},\partial_{2}\}=\left\{\frac{\partial}{\partial\mu},\frac{\partial}{\partial\sigma}\right\}$ be the basis associated with the $\theta$-coordinates, and $\{\partial_{\xi_{1}},\partial_{\xi_{2}}\}$ the basis in the $\xi$-coordinate system.
+Using the chain rule and the inverse Jacobian, the basis transformation is
+$\partial_{\xi_{\alpha}}=\sum\limits_{i=1}^{2}\bar{B_{\alpha}}^{i}\partial_{i}=\bar{B_{\alpha}}^{1}\partial_{1}+\bar{B_{\alpha}}^{2}\partial_{2},\;and\;\partial_{\xi_{\beta}}=\sum\limits_{j=1}^{2}\bar{B_{\beta}}^{j}\partial_{j}=\bar{B_{\beta}}^{1}\partial_{1}+\bar{B_{\beta}}^{2}\partial_{2}.$
+Then, the components of the Riemannian metric in the $\xi$-coordinates are given by the inner product
 $g_{\alpha\beta}(\xi)=<\partial_{\xi_{\alpha}},\partial_{\xi_{\beta}}>,\;\alpha,\;\beta\in\left\{1,2\right\}.$
-or $\bar{B_{\alpha}}^{i}=\left(\frac{\partial\theta_{i}}{\partial\xi_{\alpha}}\right)$ with
-$\partial_{\xi_{\alpha}}=\sum\limits_{i=1}^{2}\bar{B_{\alpha}}^{i}\partial_{i}=B_{\alpha}^{1}\partial_{1}+B_{\alpha}^{2}\partial_{2},\;and\;\partial_{\xi_{\beta}}=\sum\limits_{j=1}^{2}\bar{B_{\beta}}^{j}\partial_{j}=B_{\beta}^{1}\partial_{1}+B_{\beta}^{2}\partial_{2}.$
-We have
-$g_{\alpha\beta}(\xi)$ $=$ $<B_{\alpha}^{1}\partial_{1}+B_{\alpha}^{2}\partial_{2},B_{\beta}^{1}\partial_{1}+B_{\beta}^{2}\partial_{2}>$
-$=$ $<B_{\alpha}^{1}\partial_{1},B_{\alpha}^{1}\partial_{1}>+<B_{\alpha}^{1}\partial_{1},B_{\alpha}^{2}\partial_{2}>+<B_{\alpha}^{2}\partial_{2},B_{\alpha}^{1}\partial_{1}>+<B_{\alpha}^{2}\partial_{2},B_{\alpha}^{2}\partial_{2}>$
+or $\bar{B_{\alpha}}^{i}=\left(\frac{\partial\theta_{i}}{\partial\xi_{\alpha}}\right)$ We have
+$g_{\alpha\beta}(\xi)$ $=$ $<\bar{B_{\alpha}}^{1}\partial_{1}+\bar{B_{\alpha}}^{2}\partial_{2},\bar{B_{\beta}}^{1}\partial_{1}+\bar{B_{\beta}}^{2}\partial_{2}>$
+$=$ $<\bar{B_{\alpha}}^{1}\partial_{1},\bar{B_{\beta}}^{1}\partial_{1}>+<\bar{B_{\alpha}}^{1}\partial_{1},\bar{B_{\beta}}^{2}\partial_{2}>+<\bar{B_{\alpha}}^{2}\partial_{2},\bar{B_{\beta}}^{1}\partial_{1}>+<\bar{B_{\alpha}}^{2}\partial_{2},\bar{B_{\beta}}^{2}\partial_{2}>$
 $=$ $\bar{B_{\alpha}}^{1}\bar{B_{\beta}}^{1}g_{11}(\theta)+\bar{B_{\alpha}}^{2}\bar{B_{\beta}}^{2}g_{22}(\theta)$
-$\forall\alpha,\;\beta\in\left\{1,2\right\}$. ∎
+$\forall\alpha,\;\beta\in\left\{1,2\right\}$. So,
+$g_{\alpha\beta}(\xi)=\frac{1}{\sigma^{2}}\bar{B_{\alpha}}^{1}\bar{B_{\beta}}^{1}+\frac{2}{\sigma^{2}}\bar{B_{\alpha}}^{2}\bar{B_{\beta}}^{2}$
 While being (12) and (2) there is the proof of following theorem
 Theorem 3.
-Let $\xi=\left(\mu,\mu^{2}+\sigma^{2}\right)$ the dual coordinate system of $\theta-$ coordinate system on gaussian manifold. The riemannian metric $G$ on dual coordinate system $\xi$ is given by
+Let $\xi=\left(\mu,\mu^{2}+\sigma^{2}\right)$ be the dual coordinate system associated with the natural coordinates $\theta=\left(\mu,\sigma\right)$ on the Gaussian statistical manifold.
+Then, the Riemannian metric $G_{d}$ in the $\xi$-coordinate system is given by:
 $G_{d}=\frac{1}{\sigma^{4}}\left(\begin{array}[]{cc}\sigma^{2}+2\mu^{2}&-\mu\\ -\mu&\frac{1}{2}\end{array}\right),$
-the determinant of this matrix is given by
-$DetG_{d}=\frac{1}{2\sigma^{2}}$
-In a riemannian manifold, the matrix of information of Fisher in the frame of reference $\xi$ noted $g_{\alpha\beta}$ is given by
+The determinant of this metric tensor is:
+$DetG_{d}=\frac{1}{2\sigma^{6}}$
+The inverse of the Fisher information matrix in the $\xi$-coordinate system, denoted $G_{d}^{-1}$, is given by:
 $G_{d}^{-1}=\left(\begin{array}[]{cc}\sigma^{2}&2\mu\sigma^{2}\\ 2\mu\sigma^{2}&2\sigma^{4}+4\mu^{2}\sigma^{2}\end{array}\right)$
@@
 $=$ $\frac{-\mu}{\sigma^{4}}$
-We have the result ∎
+we have the following metric
+$G_{d}=\frac{1}{\sigma^{4}}\left(\begin{array}[]{cc}\sigma^{2}+2\mu^{2}&-\mu\\ -\mu&\frac{1}{2}\end{array}\right),$
+and
+$DetG_{d}=\frac{1}{2\sigma^{6}}$
+So, we have the result ∎
 4 Torsion and curvature tensor in $\xi$-coordinates system.
 Theorem 4.
-Let $\xi=\left(\mu,\mu^{2}+\sigma^{2}\right)$ the dual coordinate system of $\theta-$ coordinate system on gaussian manifold. The torsion tensor is not free.
+Let $\nabla^{(1)}$ be the $\alpha$-connection of Amari with $\alpha=1$.
+Then, the torsion tensor $T^{\gamma}_{\alpha\beta}(\xi)$ in the dual coordinate system $\xi=\left(\mu,\mu^{2}+\sigma^{2}\right)$ is non-zero.
+This is in contrast to the natural coordinate system $\theta$, where the torsion vanishes for the Levi-Civita connection.
+The presence of torsion in $\xi$-coordinates reflects the fact that duality does not preserve symmetry of the connection.
 Proof.
@@
 Theorem 5.
-The curvature tensor in $\xi$- coordinates system of gaussian distribution is given by
+Let $\xi=\left(\mu,\mu^{2}+\sigma^{2}\right)$ be the dual coordinate system associated with the Gaussian statistical manifold.
+Then, the components of the Riemann curvature tensor of the affine connection $\nabla^{(1)}$ in these coordinates include:
 $R_{1111}(\xi)$ $=$ $R_{2222}(\xi)=R_{1122}(\xi)=R_{1112}(\xi)=R_{2121}(\xi)=R_{2211}(\xi)=0$
@@
 $\Gamma_{22}^{2}(\xi)$ $=$ $\frac{1}{2}g^{21}(\xi)\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}+\frac{1}{2}g^{22}(\xi)\frac{\partial g_{22}(\xi)}{\partial\xi_{2}}$
-$\Gamma_{22}^{2}(\xi)$ $=$ $0$
+$\Gamma_{22}^{2}(\xi)$ $=$ $0.$
 $\Gamma_{12}^{2}(\xi)$ $=$ $\frac{1}{2}g^{21}(\xi)\left(\frac{\partial g_{21}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}\right)+\frac{1}{2}g^{22}(\xi)\left(\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}\right)$
@@
 $\Gamma_{22}^{1}(\xi)$ $=$ $\frac{1}{2}g^{11}(\xi)\left(\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}+\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}\right)+\frac{1}{2}g^{12}(\xi)\left(\frac{\partial g_{22}(\xi)}{\partial\xi_{2}}+\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}\right)$
-$\Gamma_{22}^{1}(\xi)$ $=$ $0$
+$\Gamma_{22}^{1}(\xi)$ $=$ $0.$
 $\Gamma_{12}^{1}(\xi)$ $=$ $\frac{1}{2}g^{11}(\xi)\left(\frac{\partial g_{21}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{11}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{12}(\xi)}{\partial\xi_{1}}\right)+\frac{1}{2}g^{12}(\xi)\left(\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{11}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{1^{\prime}2^{\prime}}}{\partial\xi^{1}}\right)$
@@
 $-<\Gamma_{11}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{1}}>-<\Gamma_{11}^{2}\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{1}}),\partial_{\xi_{1}}>$
-$=$ 00 0
+$=$ $0$
 $R_{2222}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{22}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{22}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{22}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{22}^{2}(\xi)\right)g_{22}(\xi)$
@@
 $-<\Gamma_{22}^{1}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>-<\Gamma_{22}^{2}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
-$=$ 00 0
+$=$ $0$
 $R_{1122}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{12}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{12}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{12}^{2}(\xi)-\partial_{1^{\prime}}\Gamma_{12}^{2}(\xi)\right)g_{22}(\xi)$
@@
 $-<\Gamma_{12}^{1}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>-<\Gamma_{12}^{2}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
-$=$ $0$
+$=$ $0.$
 $R_{1112}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{11}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{11}^{2}(\xi)-\partial_{1^{\prime}}\Gamma_{11}^{2}(\xi)\right)g_{22}(\xi)$
@@
 $-<\Gamma_{11}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>-<\Gamma_{11}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
-$=$ 00 0
+$=$ $0$
 $R_{2121}(\xi)$ $=$ $\left(\partial_{2^{\prime}}\Gamma_{12}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{22}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{12}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{22}^{2}(\xi)\right)g_{21}(\xi)$
@@
 $-<\Gamma_{22}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{1}}>-<\Gamma_{22}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{1}}>$
-$=$ 00 0
+$=$ $0$
 $R_{2211}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{21}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{21}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{21}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{21}^{2}(\xi)\right)g_{21}(\xi)$
@@
 $+\Gamma_{211}(\xi)\Gamma_{21}^{1}(\xi)+\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{211}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)$
-$=$ $0$
+$=$ $0.$
 $R_{2221}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{22}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{22}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{22}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{22}^{2}(\xi)\right)g_{21}(\xi)$
@@
 $+\Gamma_{211}(\xi)\Gamma_{22}^{1}(\xi)+\Gamma_{221}\Gamma_{22}^{2}(\xi)-\Gamma_{211}(\xi)\Gamma_{22}^{1}(\xi)-\Gamma_{221}(\xi)\Gamma_{22}^{2}(\xi)$
-$=$ $0$
+$=$ $0.$
 $R_{2212}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{21}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{21}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{21}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{21}^{2}(\xi)\right)g_{22}(\xi)$
@@
 $+\Gamma_{211}\Gamma_{21}^{1}(\xi)+\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{211}\Gamma_{21}^{1}(\xi)-\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)$
-$=$ $0$
+$=$ $0.$
 $R_{2122}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{12}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{22}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{12}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{22}^{2}(\xi)\right)g_{22}(\xi)$
@@
 $+\Gamma_{212}(\xi)\Gamma_{12}^{1}(\xi)+\Gamma_{222}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{211}\Gamma_{22}^{1}(\xi)-\Gamma_{122}(\xi)\Gamma_{22}^{2}(\xi)$
-$=$ $0$
+$=$ $0.$
 $R_{1221}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{22}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{12}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{22}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{12}^{2}(\xi)\right)g_{21}(\xi)$
@@
 $+\Gamma_{111}(\xi)\Gamma_{21}^{1}(\xi)+\Gamma_{121}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{221}(\xi)\Gamma_{11}^{2}(\xi)-\Gamma_{211}(\xi)\Gamma_{11}^{1}(\xi)$
-$=$ $0$
+$=$ $0.$
 $R_{1121}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{12}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{12}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{12}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{12}^{2}(\xi)\right)g_{21}(\xi)$
@@
 $+\Gamma_{111}(\xi)\Gamma_{12}^{1}(\xi)+\Gamma_{121}(\xi)\Gamma_{12}^{2}(\xi)-\Gamma_{111}(\xi)\Gamma_{12}^{1}(\xi)-\Gamma_{121}(\xi)\Gamma_{12}^{2}(\xi)$
-$=$ $0$
+$=$ $0.$
 $R_{2111}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{11}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{11}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{2}(\xi)\right)g_{21}(\xi)$
@@
 5 General conclusion
-In conclusion, on a Gaussian manifold precisely in the dual coordinate system the torsion and scalar curvature is non-zero.
-The notion of duality does not preserve the non-linear properties which are curvature and torsion in the $\theta-$coordinate system.
+This study has examined the differential geometric structure of the Gaussian statistical model when expressed in a dual coordinate system $\xi=(\mu,\mu^{2}+\sigma^{2})$.
+Through explicit computations of the Fisher information metric, the affine connection, and the curvature tensors, we have shown that the dual parametrization introduces non-trivial geometric effects.
+In contrast to the classical coordinate system $\theta=(\mu,\sigma)$, where the manifold is known to be both flat and torsion-free, the $\xi$-system exhibits non zero torsion and scalar curvature.
+These findings highlight an important phenomenon in information geometry: duality does not guarantee invariance of geometric quantities such as curvature and torsion.
+In particular, our results show that the choice of coordinate system can deeply influence the global geometric interpretation of a statistical model, even for well understood cases like the normal distribution.
+From a broader perspective, this work invites further exploration into the stability and variability of geometric invariants under coordinate transformations.
+Understanding which geometric properties are intrinsic to the model and which depend on parametrization remains a central question.
+Future research could aim to classify transformations that preserve curvature or to extend the analysis to other statistical families, such as the beta, Weibull, or log-normal distributions.
+Ultimately, this study reinforces the idea that the geometry of statistical models is not only local and computational it is also deeply sensitive to the way we represent information.
 […]